### v1

# Temporally Compressed 3D Gaussian Splatting for Dynamic Scenes

Abstract
footnotetext: Equal Contribution.
Recent advancements in high-fidelity dynamic scene reconstruction have leveraged dynamic 3D Gaussians and 4D Gaussian Splatting for realistic scene representation. However, to make these methods viable for real-time applications such as AR/VR, gaming, and rendering on low-power devices, substantial reductions in memory usage and improvements in rendering efficiency are required. While many state-of-the-art methods prioritize lightweight implementations, they struggle in handling scenes with complex motions or long sequences. In this work, we introduce Temporally Compressed 3D Gaussian Splatting (TC3DGS), a novel technique designed specifically to effectively compress dynamic 3D Gaussian representations. TC3DGS selectively prunes Gaussians based on their temporal relevance and employs gradient-aware mixed-precision quantization to dynamically compress Gaussian parameters. It additionally relies on a variation of the Ramer-Douglas-Peucker algorithm in a post-processing step to further reduce storage by interpolating Gaussian trajectories across frames. Our experiments across multiple datasets demonstrate that TC3DGS achieves up to 67$\times$ compression with minimal or no degradation in visual quality.
Figure 1: Our method, TC3DGS, achieves up to 67× compression with minimal loss in image quality, while demonstrating significantly higher rendering speed compared to [26].
1 Introduction
Dynamic scene reconstruction is essential for applications in virtual and augmented reality, gaming and robotics, where a real-time and accurate representation of moving objects and their environment is key to immersive experiences. Recent advancements such as Neural Radiance Fields (NeRF) [28] have enabled high-fidelity scene generation, albeit at the cost of slow rendering speeds. To address this limitation, 3D Gaussian Splatting (3DGS) [20] leverages sparse Gaussian splats for efficient scene rendering, particularly for static scenes.
Since the advent of 3DGS, a plethora of extensions to dynamic scenes have been proposed [39, 26, 6, 40, 41, 8]. In this context, some methods [39, 26, 6, 40] allow the Gaussians to evolve over time, capturing time-varying properties such as position, opacity, and covariance. In contrast, other methods learn spatio-temporal Gaussians to represent the dynamic content directly, which allows for a more flexible modeling of scene variations [6, 40, 8, 41]. However, obtaining a high-quality representation with these methods often requires a large number of Gaussians, leading to substantial storage overhead. Furthermore, as shown in Figure 2, spatio-temporal methods face the challenge of effectively adjusting the Gaussian parameters when applied to dynamic scenes with rapid and complex motions, such as those in [18]. This is alleviated by Dynamic 3D Gaussians [26], which enforces consistency across all frames, but at the cost of increased storage and rendering time.
To address these challenges, we propose Temporally Compressed 3D Gaussian Splatting (TC3DGS), a novel approach designed to efficiently compress dynamic 3D Gaussian representations for high-quality, real-time scene rendering. Unlike traditional methods, TC3DGS reduces both the number and the memory footprint of the Gaussians, by selectively pruning splats based on temporal importance and learning a parameter-specific bit-precision. This selective compression allows us to maintain scene fidelity while significantly reducing storage and computational requirements, making TC3DGS well-suited for dynamic environments with complex motions.
Our approach begins with a pruning and masking strategy designed to eliminate redundant Gaussians. While previous works [42, 11, 21] have proposed pruning methods for 3DGS, they do not attempt to model dynamic scenes and thus take no advantage of temporal compression. Unfortunately, adapting these pruning strategies to dynamic scenes is not straightforward, as different Gaussians may need to be pruned across different frames. Here, we introduce a method that explicitly integrates this temporal aspect into the training objective, allowing us to prune dynamic Gaussian splats more effectively. To further optimize memory usage in our scene representation, we also aim to compress the storage size of the remaining Gaussians. To this end, we developed a gradient-aware mixed-precision quantization method that adjusts the bit precision of each Gaussian parameter based on its sensitivity. We use gradient magnitudes to determine parameters with high sensitivity, and allocate them a higher precision, while those with lower impact on the scene are quantized with fewer bits. Therefore, our method achieves a good balance between compression and reconstruction accuracy. Finally, to further enhance the efficiency of our dynamic representation, we introduce a keypoint extraction algorithm as a post-processing step. This algorithm simplifies the temporal trajectories of the Gaussian parameters, preserving only the key data points that are crucial for accurate dynamic scene representation. This greatly reduces the amount of temporal data stored, allowing for more compact scene representations, as demonstrated in our experiments where we achieve a compression rate of up to 67$\times$ while preserving rendering quality.
Our experiments on benchmark datasets demonstrate the benefits of our method across diverse scenarios. Furthermore, we perform an ablation study to demonstrate the contribution of each key component in our pipeline. In a nutshell, our major contributions are as follows:
  • We introduce the first method to prune dynamic Gaussian splats, going beyond previous pruning techniques, designed for static scenes, by incorporating temporal relevance into the pruning process.
  • We develop a sensitivity-driven, gradient-based quantization method that dynamically assigns bit precision to parameters based on their impact on reconstruction accuracy, optimizing memory usage.
  • We propose a keypoint extraction post-processing algorithm to further reduce storage requirements by simplifying the Gaussian parameter trajectories, retaining only essential data points for compact scene representations.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth Dynamic 3DGS [26] TC3DGS (Ours) STG [23] 4D Gaussian [38]
Figure 2: Comparative Evaluation on Panoptic Dataset [18] Recent lightweight state-of-the-art methods (STG [23], and 4D-Gaussian [38]) struggle to accurately reconstruct scenes in complex environments. In contrast, our compression strategy effectively captures and maintains high fidelity in dynamic scene details.
2 Related Work
Dynamic 3D reconstruction.
Recent advancements in 3D reconstruction based on Neural Radiance Fields (NeRFs) [28] and 3D Gaussian Splatting (3DGS) [20] have achieved remarkable levels of visual fidelity and accuracy. These methods have subsequently been extended to 4D representations [33, 24, 13], enabling dynamic scene reconstruction. Methods to decompose a 4D scene into multiple 2D planes to learn a more compact representation are also explored by various methods [16, 3, 35, 1]. In the case of 3DGS, where Gaussians are explicitly stored and rendered, different approaches have emerged to model their time dependence. One prominent line of work [39, 26, 6, 40] in this area optimizes a canonical set of Gaussians from the initial frame, and combines it with a deformation motion field allowing temporal variations of the Gaussian parameters. However, these methods are limited to short videos, as they cannot add Gaussians after the initial frame.
Another class of methods [41, 8, 19, 9] directly models temporal Gaussians that can be present in a subset of frames, enabling certain elements to appear in selected time ranges and thus increasing the expressivity of their reconstruction. However, a major limitation across these methods is that both training and inference times for novel view synthesis scale with the number of Gaussians, the length of the sequence, and the complexity of their parameters, presenting a key bottleneck in enhancing reconstruction quality.
Compressed 3D radiance fields.
An important research direction has thus emerged in developing more compact representations of radiance fields. For NeRFs, compact grid structures [30, 4, 15, 14] have already proven effective in reducing network sizes and enhancing accuracy.
With 3DGS, recent works have either concentrated on optimizing the representation of the Gaussian parameters [11], or on identifying low-importance Gaussians and pruning them entirely [42, 11, 21]. Unorthodox techniques like representing the Gaussian parameters as 2D grids and applying image compression techniques [29, 31] have also been studied. Better initialization and weighted sampling based pruning [12] has also shown promising results. Additionally, the use of traditional compression techniques such as vector quantization [21, 31] or entropy models [5] have shown some potential for the compression of static scenes, but scaling them to dynamic scenes with possibly hundreds or thousands of frames remains a challenge.
Indeed, dynamic scenes require an even larger set of parameters to accurately capture motion, temporal variations, and complex interactions within the scene.
The temporal relevance of each Gaussian changes dynamically, and traditional pruning strategies designed for static scenes are insufficient, as they lack adaptability to these fluctuations. To the best of our knowledge, we are the first to propose a compression framework specifically designed for dynamic 3D Gaussians. Our approach combines temporal relevance-based pruning, gradient-based mixed-precision quantization, and trajectory simplification to address the unique requirements of dynamic scenes.
3 Method
In this section, we will first briefly discuss the Dynamic 3DGS [26] method in Sec. 3.1, which serves as the foundation of our proposed approach. We then detail in Sec. 3.2.1 our novel masking and pruning strategies designed to eliminate redundant Gaussians. Next, we present a sensitivity-based mixed precision technique for efficient parameter compression in Sec. 3.2.2, followed by a post-training compression strategy aimed at minimizing storage overhead in Sec. 3.2.3.
Figure 3: Overview of our Temporally Compressed 3D Gaussian Splatting for Dynamic Scenese (TC3DGS) method. Our approach involves a temporally consistent masking strategy to select relevant 3D Gaussians across frames. The masked Gaussians are then pruned and quantized using a gradient-based, parameter-aware bit-precision quantization scheme, which optimizes memory usage while preserving scene fidelity.
3.1 Dynamic 3D Gaussians
Dynamic 3DGS [26] models dynamic scenes by allowing the Gaussians to move and rotate over time while enforcing that they have persistent color, opacity, and scale. This approach reconstructs a dynamic 3D scene over time using a series of images taken from multiple cameras across different time steps, along with the cameras’ intrinsic and extrinsic parameters. The method sequentially reconstructs a dynamic 3D scene by initializing each time step from the previous one and performing test-time optimization without additional training data. This is unlike most other concurrent methods, which attempt to optimize all frames jointly. This conceptual difference allows Dynamic 3DGS [26] to enforce consistency in the reconstructed Gaussians across frames. Additionally, this method enables tracking objects throughout the dynamic scene, as each 3D Gaussian has a unique correspondence across frames.
Each dynamic scene $\mathcal{S}$ is parameterized by a set of Dynamic 3D Gaussians, with certain parameters fixed for the entire sequence and others varying over time. Specifically, each Gaussian retains a fixed 3D scale $(s_{x},s_{y},s_{z})$, color $(r,g,b)$, opacity logit $o$, and background logit $bg$ as determined in the initial frame. In contrast, parameters such as the 3D center $(x_{t},y_{t},z_{t})$ and 3D rotation, expressed as a quaternion $(q_{w,t},q_{x,t},q_{y,t},q_{z,t})$, evolve over time. This formulation enables the Gaussians to represent consistent scene elements across frames while dynamically adjusting their positions and orientations.
During training, this 3D scene representation is iteratively updated using a differentiable renderer to minimize the photometric error between the rendered and input images. Specifically, each Gaussian influences a point $p$ in physical 3D space according to the standard (unnormalized) Gaussian equation weighted by its opacity, i.e.,
$f_{i,t}(p)$ $=\sigma({o_{i}})\exp\left(-\frac{1}{2}(p-\mu_{i,t})^{T}\Sigma_{i,t}^{-1}(p-\mu_{i,t})\right)\;,$
[…]
3.2 TC3DGS
Dynamic 3DGS [26] is particularly promising because it models the dynamic scene as movements of Gaussians under kinematic constraints w.r.t. the previous timestep. By optimizing the position and rotation of the Gaussians instead of learning deformation functions, it removes the limitation on possible deformations due to the characteristics of the modeling function. However, by learning position and rotation at each timestep separately, the number of parameters increases linearly, resulting in large storage sizes, and thus limiting its applicability to high-fidelity and long-range scene modeling.
3.2.1 Gaussian Masking and Pruning
Pruning techniques aimed at identifying and removing low-importance Gaussians have been successfully applied to static 3D Gaussian splatting [21, 32, 17, 11]. However, for dynamic scenes, a temporally consistent importance measure is required to ensure that the pruned Gaussians remain insignificant throughout the scene’s duration.
Various approaches to computing the importance of each Gaussian in static scenes using training images and camera positions have bee proposed [32, 11]. These methods focus on the contribution of each Gaussian to the training views. In dynamic scenes, Gaussians are not stationary, so their contributions vary over time. To prune Gaussians effectively from dynamic scenes, it is essential to maintain the contributions of high-importance Gaussians consistently high, while suppressing low-importance ones, thus enabling more effective pruning.
Compact-3DGS [21] introduces a masking approach based on Gaussian volume and transparency. Gaussians with low opacity, minimal volume, or both are masked out since they have negligible impact on the rendered images. A mask parameter $m\in\mathbb{R}^{N}$ is learned to produce binary masks $M\in\{0,1\}^{N}$ using a straight-through estimator. This binary mask is then applied to the Gaussians by scaling their opacities and sizes. This is expressed as
$M_{n}$ $=\operatorname*{sg}(\mathds{1}[\sigma(m_{n})>\epsilon]-\sigma(m_{n}))+\sigma(m_{n}),$ (1)
$\hat{s}_{n}$ $=M_{n}s_{n},\quad\hat{o}_{n}=M_{n}o_{n},$ (2)
where $n$ represents the Gaussian index, $\epsilon$ denotes the masking threshold, $\operatorname*{sg}(\cdot)$ is the stop-gradient operator, and $\mathds{1}[\cdot]$ and $\sigma(\cdot)$ correspond to the indicator and sigmoid functions, respectively.
Compact-3DGS [21] explains that this mask learns to remove Gaussians with low opacity and/or small volume. However, as shown in Figure 4, this explanation does not hold in dynamic scenes. The value of $m_{n}$ can change significantly across time frames, even with fixed opacity $o$ and scale $s$ after the initial time frame.
Figure 4: Mask Consistency. (Left) Mask values of Gaussians trained independently for each time frame. (Right) Mask values trained with our loss.
In dynamic scenes, $m_{n}$ depends on the 2D projected area of the Gaussians in the training views and their transmittance, where transmittance $T_{i}$ for the $i^{th}$ Gaussian along a camera ray is defined as the Gaussian’s contribution to blending, i.e.,
$T_{i}=\sigma(o_{i})\prod_{j=1}^{i-1}(1-\sigma(o_{j})).$ (3)
As a Gaussian moves relative to the others, changes in $T_{i}$ are reflected in $m_{n}$. Similarly, when a Gaussian moves toward or away from the training cameras, its projected area in the views changes, affecting $m_{n}$ accordingly.
Since the values in $m$ strictly need to be high due to their role in rendering through Eq. 1 and the photometric loss, we introduce an additional regularization to incentivize lower values. We thus regularize $m$ by minimizing
$\mathcal{L}_{mask}=\sum_{n=1}^{N}\sigma(m_{n}).$ (4)
This regularization loss penalizes unnecessarily high values, reducing $m$ to the minimum required to produce satisfying renderings. The flexibility of this learned mask is one of its key advantages, as it can be optimized to exhibit desired properties via the use of additional constraints.
A key motivation for ensuring consistency of $m_{n,t}$ across frames is to capture the global importance of the Gaussians during pruning. To achieve this, we learn $m_{n,t}$ for each timestamp and introduce a consistency loss function that encourages $m_{n,t}$ to remain close to $m_{n,t-1}$. Specifically, we define this mask consistency loss function as
$\mathcal{L}_{mc}=\sum_{n=1}^{N}|m_{n,t}-sg(m_{n,t-1})|.$ (5)
It ensures that the masks exhibit stability across the frames, as shown in Figure 4 (Right). By maintaining consistency, our approach prevents sudden fluctuations in Gaussian importance, which can degrade rendering quality and lead to suboptimal pruning results.
After optimizing across all timestamps, we perform Gaussian pruning based on the average value of $m_{n}$. Specifically, a Gaussian is pruned if its average value across all timestamps satisfies
$\frac{1}{T}\sum_{t=1}^{T}\sigma(m_{n,t})<\epsilon.$ (6)
3.2.2 Gradient-Aware Mix-Precision Quantization
The influence of the Gaussian parameters on reconstruction quality is highly variable; a small adjustment in one parameter can significantly alter the rendered image, whereas similar adjustments in other parameters or in the same parameter of a different Gaussian may have minimal impact. Our approach uses gradient-based sensitivities to dynamically assign bit precision to each parameter, based on its influence on reconstruction accuracy. By leveraging this adaptive, in-optimization quantization, each parameter adjusts its quantization scale [10] in real-time, preserving detail in the reconstructed scene.
We first calculate the mean sensitivity for each parameter based on the gradients [43, 32], reflecting each parameter’s impact on the image reconstruction performance. We introduce a sensitivity coefficient $S(\theta)$, which represents the responsiveness of image quality $Q$ to changes in a Gaussian parameter $\theta$. This coefficient is formulated as
$S(\theta)$ $=\frac{1}{\sum_{k=1}^{K}N_{k}}\sum_{k=1}^{K}\big{|}\frac{\partial Q_{k}}{\partial\theta}\big{|}.$ (7)
where $K$ is the total number of training images used for reconstruction, $N_{k}$ is the number of pixels in the $k^{th}$ image, and $Q_{m}$ denotes the cumulative pixel intensity across the RGB channels in image $k$, serving as a proxy for image quality.
The coefficient $S(\theta)$ quantifies how sensitive $Q$ is to variations in $\theta$, measured by the gradient $\frac{\partial Q}{\partial\theta}$. A higher gradient indicates that small adjustments in $\theta$ yield large changes in image quality, suggesting a greater impact of $\theta$ on the reconstruction.
By using this impact-based metric, we can effectively rank the parameters by their importance on image fidelity. We then normalize each sensitivity coefficient by scaling it based on the minimum and maximum co-efficient across all parameters. This standardization ensures that all the coefficients fall within a consistent range.
Using the normalized sensitivity $\gamma\in[0,1]$, we assign a dynamic bit precision $b$ within a range of bits $b\in[b_{\text{min}},b_{\text{max}}]$, where $b\in\mathbb{Z}$ is calculated as
$b=\left\lfloor\gamma\cdot(b_{\text{max}}-b_{\text{min}})\right\rfloor+b_{\text{min}}.$ (8)
Figure 5: Keypoint Interpolation. In this example, we represent a position across 150 frames with only 5, 4 and 6 keypoints for $x$, $y$ and $z$, respectively, with only $0.038$ MSE. By comparison, uniformly sampling 7 keypoints increases storage and increases error to $0.089$ MSE.
This approach allocates higher bit precision to more sensitive parameters and lower bit precision to less sensitive ones, optimizing the balance between computational efficiency and model accuracy.
After running our scene reconstruction process for a specified number of iterations, we apply mixed-precision quantization to all Gaussian parameters, excluding the position parameter, to achieve low-bit precision for the other parameters. Instead of relying on traditional vector quantization (VQ) or basic min-max quantization, we propose a parameter quantization technique with learnable scaling factors [10], integrating it directly into the optimization process rather than treating it as a post-optimization fine-tuning step. Consequently, many parameters can be effectively quantized to even 4-bit precision, reducing memory and computational load without compromising reconstruction quality.
Training. We train the Gaussian parameters to model the scene one time frame at a time. Following [26], we use physically-based priors to regularize the Gaussians. The optimization objective is defined as
$\mathcal{L}=\mathcal{L}_{original}+\lambda_{mask}\mathcal{L}_{mask}+\lambda_{mc}\mathcal{L}_{mc},$ (9)
where $\mathcal{L}_{original}$ is the same loss function as in [26], while $\mathcal{L}_{mask}$ and $\mathcal{L}_{mc}$ are defined in Eq. 4 and 5, respectively.
Algorithm 1 Keypoint Selection
Input: values, max_keypoints ($max_{kp}$), tolerance ($\tau$)
1:Initialize keypoints at first and last positions
2:for $i=1$ to $max_{kp}-2$ do
3:     Compute interpolated values based on keypoints
4:     Compute error for every value
5:     Compute error (mse) for entire sequence
[…]
7:         break
8:     end if
9:     Select the value with the highest error and add it to keypoints
10:end for
3.2.3 Keypoint Interpolation
Previous implementations of dynamic 3DGS often model the time-dependence of parameters by fitting polynomials to the Gaussian parameters, which greatly limits the range of motions that can accurately be modeled. By contrast, Dynamic 3DGS [26] opts for inefficiently storing all time-dependent parameters, such as Gaussian means, rotations, and colors, for all time frames. While this greatly increases the expressivity of this method compared to other works, it comes at a significant cost in memory.
We take a different approach and observe that only a small subset of keypoints are required to accurately reconstruct complex motions. However, the placement of these keypoints across time cannot be predetermined, as it depends on the individual Gaussians. For instance, background Gaussians require a single keypoint to cover the whole sequence, whereas moving objects will need substantially more keypoints. This motivates the development of our keypoint selection strategy, which we adapt from the Ramer–Douglas–Peucker (RDP) algorithm [7]. It is applied as a post-processing step to further reduce storage requirements for the Gaussian parameters that change over time in dynamic scenes.
While the RDP algorithm selects keypoints from a sequence based on a local error tolerance, $\xi$, we propose a novel keypoint selection method, as outlined in Algorithm 1 . This method offers several advantages over the RDP algorithm. It provides greater flexibility by allowing control via both an acceptable tolerance value, $\tau$, and a maximum number of keypoints. The parameter $\tau$ defines the maximum allowable Mean Squared Error (MSE) over the sequence. Unlike RDP, which is solely controlled by $\xi$, our method enables a hard maximum bound on the number of keypoints, allowing for more precise, fine-grained control.
Following this, we flatten and transpose the time-dependent parameters from $\mathbb{R}^{T\times N\times D}$ to $\mathbb{R}^{ND\times T}$, transforming the data into $ND$ sequences of length $T$. We then compute the keypoints for all Gaussians in parallel, forming sparse matrices. The sparsity of these matrices is controlled using the parameters $max_{kp}$ and $\tau$, and and we store these sparse matrices to minimize memory usage. Upon loading the scene data, we interpolate the keypoints across all timesteps to reconstruct the dense matrices, which are then reshaped to their original parameter dimensions.
4 Experiments
We compare our method with multiple techniques extending the 3D Gaussians approach to dynamic scenes. We evaluate the methods on diverse datasets covering different real-world and synthetic scenarios. Results on additional datasets are provided in the supplementary material along with more visualizations.
4.1 Datasets
4.1.1 Panoptic Sports Dataset
We evaluate our method on the Panoptic Sports dataset, a subset from the Panoptic Studio dataset [18]. This dataset contains 6 different scenes, each having 31 camera sequences spanning 150 frames. We use 4 cameras (0, 10, 15 and 30) for testing and the rest for training, following the convention set by [26]. In addition to the images, we use the provided foreground/background segmentation to apply a segmentation loss to improve the results and prevent the background from moving. This dataset contains complex motions with objects moving quickly and over long distances.
Table 1: Quantitative results on the Panoptic dataset. The best result is shown in bold, while the second-best result is underlined. The rendering resolution is set to 640$\times$460.
Model PSNR (dB)↑ SSIM↑ LPIPS↓ FPS ↑ Storage (MB)↓
Dynamic 3DGS [26] 28.7 0.91 0.17 760 1994
STG [23] 20.5 0.78 0.44 429 19
4D Gaussian [38] 27.2 0.91 0.18 40 62
TC3DGS (Ours) 28.3 0.91 0.17 890 49
4.1.2 Neural 3D Video Dataset
The Neural 3D Video dataset [22] consists of 6 scenes with a number of cameras ranging from 18 to 21 and sequences of 300 frames. The images in this dataset are or resolution 2704$\times$2028, but we downsample them to 1352$\times$1014 for fair comparison with other methods. We hold out camera 0 for testing and do not apply any background loss, as background segmentations are not available for this dataset.
4.2 Implementation Details
For our experiments, we closely follow the hyperparameters outlined in Dynamic 3DGS [26]. Specifically, for the compression strategy, we set both the mask weight parameter $\lambda_{\text{mask}}$ and the masking consistency loss parameter $\lambda_{\text{mask-cons}}$ to $0.01$ in our primary experiments. In terms of quantization, we use Gaussian parameters with bin sizes of $b_{\text{min}}=4$ and $b_{\text{max}}=8$, while positional data is quantized to 16-bit precision to ensure spatial accuracy. For the learnable quantization step size, we use a $0.02$ learning rate. Quantization is applied after 6000 iterations in the first scene for all experiments. For keypoint interpolation, we set the tolerance to $\tau=1\times 10^{-5}$ for the Panoptic dataset and $\tau=1\times 10^{-7}$ otherwise. Moreover, we restrict the maximum number of keypoints to $\text{max}_{\text{kp}}=30$ for the Panoptic dataset and $\text{max}_{\text{kp}}=60$ otherwise, which provides an effective balance between compression and temporal trajectory accuracy. Additional ablation studies with varying values of certain hyperparameters are provided in the supplementary material.
4.3 Results
Table 2: Quantitative comparisons with various competitive baselines on the Neural 3D Video dataset. represent methods trained with a 50-frame video sequence, requiring six models to complete the evaluation.“Storage” refers to the total model size for 300 frames.
[…]
4DGS [41] 31.57 96.69 3128.00MB
4D Gaussian [38] 31.15 30.00 90MB
C-D3DGS [19] 30.46 118.00 338.00MB
Deformable 3DGS [40] 30.98 29.62 32.64MB
E-D3DGS [2] 31.20 69.70 40.20MB
STG [23] 32.04 273.47 175.35MB
Dynamic 3DGS [26] 30.97 460.00 2772.00MB
TC3DGS (Ours) 30.58 596.32 51.34MB
(a) Ground-Truth
(b) TC3DGS Render
Figure 6: Qualitative Evaluation on the Neural 3D Video dataset. This figure shows a comparison between the ground truth and our reconstruction, demonstrating that our method achieves near-identical results (PSNR: 32.7) despite extreme compression.
Table 3: Ablation study on the proposed contributions. ‘M’, ‘Q’ and ‘I’ denote masking, quantization, interpolation, respectively. ‘#Gauss’ means the number of Gaussians.
[…]
M Q I PSNR #Gauss Storage FPS PSNR #Gauss Storage FPS
Dynamic 3DGS 28.2 349K 2161 MB 582 33.1 294K 3370 MB 472
28.1 189K 1087 MB 750 32.9 59 K 674 MB 583
27.9 189K 299 MB 750 32.8 59 K 194 MB 583
27.9 189K 46 MB 750 32.7 59 K 53 MB 583
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Dynamic 3DGS [26] Pruned Quantized + Pruned Quantized + Pruned + KPI
Figure 7: Ablation study. Adding the different components of our method comes at virtually no cost in visual accuracy but at significantly decreased storage and rendering cost.
As shown in Table 1, our method achieves results comparable to Dynamic 3DGS [26], while utilizing, on average, 40 times less storage. Similarly, Table 2 presents the results of our method on the DyNeRF dataset, demonstrating competitive performance with a significantly smaller storage footprint. Furthermore, our approach delivers the fastest rendering speed.
On the Panoptic dataset, we failed to obtain reasonable results for the STG baseline [23], while 4D Gaussian [38] produce very distorted images, as shown in Figure 2. This dataset demonstrates the strengths of explicit methods such as Dynamic 3DGS [26] and ours. While Dynamic 3DGS [26] is able to model the complex motion and perform well on the Panoptic dataset, it is highly memory inefficient. In comparison, our model performs equally well while using much less storage.
4.4 Ablation Studies
We conduct an ablation study to evaluate the effectiveness of the individual components of our method. We report the results of this ablation in Table 3 on two scenes from the Panoptic dataset [18] and the Neural 3D Video dataset [22].
These results show the importance of each component of our method in reducing the storage size of the 3DGS [26]. In these experiments, our pruning strategy divides the number of Gaussians by 2 and by 5 folds for each scene. Then, sensitivity-aware quantization reduces the necessary storage space by 5, and keypoint interpolation by an additional 4 to 5 fold. Altogether, TC3DGS yields a compression ratio of 49 and 64, respectively. Most notably, this drastic size reduction comes at little to no cost in novel view synthesis, with an insignificant drop in PSNR of at most 0.4, which is barely perceptible.
5 Limitations
Our approach has certain limitations rooted in the nature of dynamic 3D Gaussian splatting (3DGS) and the constraints of compressing temporal data for dynamic scenes. First, unlike compression approaches for static 3DGS, we are limited in our ability to aggressively compress position parameters, as doing so would compromise temporal consistency and spatial accuracy in the dynamic setting. Additionally, as our method builds on Dynamic 3DGS, it inherits its inability to accurately reconstruct new objects that enter the scene after the initial frame. This restricts its use in scenarios requiring complete adaptability to scene changes. Nonetheless, our method excels at handling scenes with complex and fast motions, maintaining high fidelity and rendering efficiency despite these constraints.
6 Conclusion
We introduced Temporally Compressed 3D Gaussian Splatting (TC3DGS), a novel framework designed to achieve memory-efficient and high-speed reconstruction of dynamic scenes. Our method achieves up to 67x compression and up to three times faster rendering while retaining high fidelity in complex motion scenarios. Through selective temporal pruning and gradient-based quantization, TC3DGS minimizes memory usage with minimal impact on visual quality. While our method is effective, limitations remain in aggressively compressing position data and handling new objects that enter the scene mid-sequence. Future work will focus on bridging the gap between spatio-temporal methods and storage-efficient dynamic scene reconstruction, aiming to extend adaptability and further reduce storage requirements for complex, evolving environments.
[…]
Supplementary Material
7 Technicolor Dataset [34]
To further demonstrate the effectiveness of our method, we evaluate our method on the Technicolor Dataset. The results in Table 4 show that our method provides competitive performance with a significantly smaller storage footprint. Figure 11 shows a comparison between the ground truth and our reconstruction, demonstrating that our method achieves near identical results.
Table 4: Per-Scene Results on the Technicolor Dataset. Our method achieves performance comparable to Dynamic3DGS [26] while significantly reducing the model size.
Method Avg. Birthday Fabien Painter Theater Train
[…]
STG [23] $33.60$ $32.09$ $35.70$ $36.44$ $30.99$ $32.58$
D. 3DGS [26] $32.12$ $30.82$ $33.62$ $34.73$ $31.12$ $30.67$
Ours $31.9$ $30.33$ $33.15$ $34.62$ $31.01$ $30.57$
[…]
STG [23] - - - - - -
D. 3DGS [26] $0.88$ $0.90$ $0.88$ $0.87$ $0.86$ $0.88$
Ours $0.85$ $0.88$ $0.86$ $0.81$ $0.84$ $0.86$
[…]
STG [23] $0.08$ $0.04$ $0.11$ $0.10$ $0.13$ $0.04$
D. 3DGS [26] $0.13$ $0.07$ $0.16$ $0.16$ $0.14$ $0.11$
Ours $0.14$ $0.09$ $0.15$ $0.17$ $0.16$ $0.13$
Table 5: Per-scene quantitative comparisons on the Panoptic Sports Dataset [18].
Method Avg. Basket. Juggle Boxes Softball Tennis Football
[…]
4DGaussians [38] $27.2$ $26.68$ $27.47$ $27.02$ $27.44$ $27.31$ $27.39$
STG [23] $20.45$ $21.60$ $19.93$ $20.65$ $19.44$ $20.82$ $20.23$
D. 3DGS [26] 28.70 28.22 29.48 29.46 28.43 28.11 28.49
Ours $27.81$ $27.92$ $29.15$ $28.28$ $27.96$ $25.97$ $28.00$
[…]
4DGaussians [38] 0.91 $0.90$ 0.92 0.91 0.92 0.92 0.92
STG [23] $0.79$ $0.78$ $0.78$ $0.79$ $0.79$ $0.78$ $0.78$
D. 3DGS [26] 0.91 0.91 0.92 0.91 $0.91$ $0.91$ $0.91$
Ours $0.89$ $0.89$ $0.90$ $0.89$ $0.89$ $0.89$ $0.89$
[…]
4DGaussians [38] $0.11$ 0.14 0.10 0.10 0.10 0.11 0.10
STG [23] 0.10 0.14 0.10 0.10 0.10 0.11 0.10
D. 3DGS [26] $0.17$ $0.18$ $0.15$ $0.17$ $0.19$ $0.18$ $0.17$
Ours $0.20$ $0.20$ $0.19$ $0.19$ $0.20$ $0.20$ $0.20$
SIZE(MB)$\downarrow$
4DGaussians [38] $63$ $66$ $59$ $63$ $57$ $72$ $65$
STG [23] 19 19 19 20 19 22 19
D. 3DGS [26] $2008$ $2161$ $1935$ $2021$ $2021$ $1915$ $2000$
Ours $49$ $46$ $32$ $44$ $41$ $48$ $34$
8 Detailed Results for Panoptic and Neural 3D Video Dataset
We present detailed results for the two datasets discussed in Sec. 4.1 of our paper, summarized in Table 5 and 8. These results include various metrics evaluated for each scene in both datasets, offering a granular view of performance. Additionally, we provide supplementary visualizations for the Neural 3D Video dataset, which further demonstrate the effectiveness of our method in capturing dynamic scenes with high fidelity. Finally, we present an ablation study over the main elements of our method in Figure 8.
9 Additional Experiments
We conducted several experiments to better understand hyperparameters of our method.
9.1 Masking and Mask Consistency
We vary the weight of the mask loss $\lambda_{\text{mask}}$ and the masking consistency loss $\lambda_{\text{mask-cons}}$ on the first 50 frames of the Basketball scene from the Panoptic Sports dataset. The results, shown in Table 6, indicate that increasing the weight of these parameters leads to greater pruning of Gaussians but results in reduced image quality. We found $\lambda_{\text{mask}}=\lambda_{\text{mask-cons}}=0.01$ to be an effective balance between maintaining image quality and minimizing model size.
(a) SSIM $\uparrow$
Mask Consistency Loss
Mask Loss 0 0.01 0.05 0.1
0.005 0.9087 0.9105 0.9109 0.9068
0.01 0.9126 0.9140 0.9113 0.9092
0.05 0.9036 0.9006 0.8986 0.8970
0.1 0.8938 0.8956 0.8949 0.8972
(b) PSNR $\uparrow$
Mask Consistency Loss
Mask Loss 0 0.01 0.05 0.1
0.005 28.0 27.8 28.2 26.8
0.01 28.1 28.3 26.8 26.9
0.05 27.2 26.6 26.0 26.0
0.1 27.1 27.4 26.6 26.7
(c) LPIPS $\downarrow$
Mask Consistency Loss
Mask Loss 0 0.01 0.05 0.1
0.005 0.1797 0.1769 0.1781 0.1815
0.01 0.1746 0.1761 0.1748 0.1829
0.05 0.2046 0.2090 0.2139 0.2172
0.1 0.2345 0.2294 0.2344 0.2335
(d) Number of Gaussians $\downarrow$
Mask Consistency Loss
Mask Loss 0 0.01 0.05 0.1
0.005 208403 155720 143112 140897
0.01 153757 112909 98017 93568
0.05 49770 47110 29629 26154
0.1 28972 31067 19472 15656
Table 6: Impact of Mask Loss and Mask Consistency Loss. Increasing the weight of these masks leads to more aggressive pruning of Gaussians, but at the cost of rendering quality. Setting $\lambda_{mask}=0.01$ and $\lambda_{mask-cons}=0.01$ provides a good trade-off between quality and storage efficiency.
9.2 Quantization
We ran experiments with sensitivity-aware quantization with different bit-ranges by varying $b_{\text{min}}$ and $b_{\text{max}}$ as well as using uniform bitwidth for all paramaters. We shown in the Table 7 that using adaptive bitwidth, the average bitwidth is lower than uniform quantization at 8 bits while image quality is similar. Whereas, compared to uniform quantization using lower bitwidth, our average bitwidth is slightly higher while image quality is improved considerably.
Bit-precision PSNR SSIM LPIPS Compression
uniform 4 bit 25.1 0.81 0.32 8x
uniform 5 bit 25.8 0.84 0.28 6.4x
uniform 6 bit 26.1 0.86 0.26 5.3x
uniform 8 bit 28.1 0.90 0.20 4x
Ours [5,8] 28.0 0.90 0.20 5.6x
Ours [4,8] 27.9 0.90 0.20 6.3x
Ours [3,8] 26.2 0.84 0.28 7.3x
Table 7: Comparison of Sensitivity-Aware and Uniform Bit Quantization. Our sensitivity-aware [4,8]-bit quantization achieves PSNR comparable to 8-bit precision while providing compression similar to uniform 5-bit quantization.
9.3 Keypoint Interpolation
We experimented with varying the hyperparameters for keypoint interpolation, specifically the maximum number of keypoints, $max_{kp}$. As shown in Figure 9, increasing $max_{kp}$ reduces compression error and improves image quality. Conversely, increasing $\tau$ allows for greater error tolerance, leading to a decrease in image quality.
The experiments summarized in Figure 9 were conducted on 150 frames of the Basketball scene with masking parameters set to $\lambda_{\text{mask}}=\lambda_{\text{mask-cons}}=0.01$ and fixed $tau=1\times e^{-6}$. We observed that overly relaxing $\tau$ leads to underfitted keypoints, resulting in compromised rendering quality. On the other hand, increasing $max_{kp}$ beyond a certain threshold results in diminishing returns, where interpolation error continues to decrease, but the improvement in image quality becomes marginal.
Additionally, Figure 10 illustrates that only the color parameters saturate the maximum keypoints, whereas positions and rotations can be adequately approximated with fewer keypoints.
We selected the hyperparameters to strike a balance between rendering quality and storage efficiency. For the Panoptic Sports dataset, we used $max_{kp}=30$ and $\tau=1\times 10^{-5}$. For the Neural 3D Video dataset, which consists of 300-frame sequences, we increased $max_{kp}$ to 60 to accommodate the longer sequence length. In the case of the Technicolor dataset, despite its long sequences, we matched the experimental setup by using 50 timesteps per scene and reduced $max_{kp}$ to 15.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Dynamic 3DGS [26] Pruned Quantized + Pruned Quantized + Pruned + KPI
Figure 8: Ablation Study on Panoptic Sports Dataset
Table 8: Per-scene quantitative comparisons on the Neural 3D Video Dataset [22]. Some methods only report part of the scenes. 1 only includes the Flame Salmon scene. 2 excludes the Coffee Martini scene. “-” denotes results that are unavailable in prior work.
Method Avg. Coffee Martini Cook Spinach Cut Roasted Beef Flame Salmon Flame Steak Sear Steak
[…]
Neural Volumes [25] 1 $22.80$ $22.80$
LLFF [27] 1 $23.24$ $23.24$
DyNeRF [22] 1 $29.58$ $29.58$
HexPlane [3] 2 $31.71$ $32.04$ $32.55$ $29.47$ $32.08$ $32.39$
NeRFPlayer [36] $30.69$ 31.53 $30.56$ $29.35$ $31.65$ $31.93$ $29.13$
HyperReel [1] $31.10$ $28.37$ $32.30$ $32.92$ $28.26$ $32.20$ $32.57$
K-Planes [16] $31.63$ $29.99$ $32.60$ $31.82$ $30.44$ $32.38$ $32.52$
MixVoxels-L [37] $31.34$ $29.63$ $32.25$ $32.40$ $29.81$ $31.83$ $32.10$
MixVoxels-X [37] $31.73$ $30.39$ $32.31$ $32.63$ 30.60 $32.10$ $32.33$
4D Gaussians [38] $30.67$ $27.34$ $32.46$ $32.90$ $29.20$ $32.51$ $32.49$
Dynamic 3DGS [26] $30.97$ $27.32$ $32.97$ $31.75$ $27.26$ $33.24$ $33.68$
STG [23] 32.05 $28.61$ 33.18 $33.52$ $29.48$ 33.64 33.89
Ours $30.58$ $27.42$ $31.92$ 33.61 $27.12$ $32.48$ $33.04$
[…]
Neural Volumes [25] 1 $0.295$ $0.295$
LLFF [27] 1 $0.235$ $0.235$
DyNeRF [22] 1 $0.083$ $0.083$
HexPlane [3] 2 $0.075$ $0.082$ $0.080$ $0.078$ $0.066$ $0.070$
NeRFPlayer [36] $0.111$ $0.085$ $0.113$ $0.144$ $0.098$ $0.088$ $0.138$
HyperReel [1] $0.096$ $0.127$ $0.089$ $0.084$ $0.136$ $0.078$ $0.077$
MixVoxels-L [37] $0.096$ $0.106$ $0.099$ $0.088$ $0.116$ $0.088$ $0.080$
MixVoxels-X [37] $0.064$ $0.081$ $0.062$ $0.057$ $0.078$ $0.051$ $0.053$
Dynamic 3DGS [26] $0.082$ $0.122$ $0.077$ $0.087$ $0.09$ $0.072$ $0.068$
STG [23] 0.044 0.069 0.037 0.036 0.063 0.029 0.030
Ours $0.089$ $0.129$ $0.081$ $0.088$ $0.094$ $0.075$ $0.071$
(a) Interpolation Error and PSNR vs Maximum Keypoints
(b) Interpolation Error and SSIM vs Maximum Keypoints
Figure 9: Impact of increasing Maximum Keypoints The figure illustrates the trend of interpolation error and PSNR/SSIM as we increase $max_{kp}$. It highlights that the rate of change in interpolation error and PSNR/SSIM diminishes as we increase $max_{kp}$
(a) Mean number of Keypoints
(b) Distribution of Keypoints
Figure 10: Distribution of Keypoints Across Different Parameters. The figure illustrates the mean number of keypoints for Gaussian parameters. It highlights that only color parameters reach $max_{kp}$ saturation, while positions and rotations can be effectively estimated with fewer keypoints.
(a) Ground-Truth
(b) TC3DGS Render
Figure 11: Qualitative Evaluation on the Technicolor dataset. This figure shows a comparison between the ground truth and our reconstruction, demonstrating that our method achieves near-identical results despite extreme compression.
(a) Ground-Truth
(b) TC3DGS Render
Figure 12: Qualitative Evaluation on the Neural 3D Video Dataset. This figure shows a comparison between the ground truth and our reconstruction, demonstrating that our method achieves near-identical results despite extreme compression. Top: Cook Spinach, Bottom: Sear Steak

### v2

@@
-Temporally Compressed 3D Gaussian Splatting for Dynamic Scenes
+Lausanne, Switzerland
 Abstract
-footnotetext: Equal Contribution.
-Recent advancements in high-fidelity dynamic scene reconstruction have leveraged dynamic 3D Gaussians and 4D Gaussian Splatting for realistic scene representation.
-However, to make these methods viable for real-time applications such as AR/VR, gaming, and rendering on low-power devices, substantial reductions in memory usage and improvements in rendering efficiency are required.
-While many state-of-the-art methods prioritize lightweight implementations, they struggle in handling scenes with complex motions or long sequences.
-In this work, we introduce Temporally Compressed 3D Gaussian Splatting (TC3DGS), a novel technique designed specifically to effectively compress dynamic 3D Gaussian representations.
-TC3DGS selectively prunes Gaussians based on their temporal relevance and employs gradient-aware mixed-precision quantization to dynamically compress Gaussian parameters.
-It additionally relies on a variation of the Ramer-Douglas-Peucker algorithm in a post-processing step to further reduce storage by interpolating Gaussian trajectories across frames.
-Our experiments across multiple datasets demonstrate that TC3DGS achieves up to 67$\times$ compression with minimal or no degradation in visual quality.
-Figure 1: Our method, TC3DGS, achieves up to 67× compression with minimal loss in image quality, while demonstrating significantly higher rendering speed compared to [26].
+[…]
+Project Page: https://ahmad-jarrar.github.io/tc-3dgs/
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Ground Truth D-3DGS [luiten2023dynamic] TC3DGS (Ours) STG [stg] 4D Gaussian [4dgs]
+Figure 1: Comparative Evaluation on Panoptic Dataset [Joo_2019] Recent lightweight state-of-the-art methods (STG [stg], and 4D-Gaussian [4dgs]) struggles to reconstruct scenes in complex environments accurately.
+In contrast, our compression strategy effectively captures and maintains high fidelity.
 1 Introduction
-Dynamic scene reconstruction is essential for applications in virtual and augmented reality, gaming and robotics, where a real-time and accurate representation of moving objects and their environment is key to immersive experiences.
-Recent advancements such as Neural Radiance Fields (NeRF) [28] have enabled high-fidelity scene generation, albeit at the cost of slow rendering speeds.
-To address this limitation, 3D Gaussian Splatting (3DGS) [20] leverages sparse Gaussian splats for efficient scene rendering, particularly for static scenes.
-Since the advent of 3DGS, a plethora of extensions to dynamic scenes have been proposed [39, 26, 6, 40, 41, 8].
-In this context, some methods [39, 26, 6, 40] allow the Gaussians to evolve over time, capturing time-varying properties such as position, opacity, and covariance.
-In contrast, other methods learn spatio-temporal Gaussians to represent the dynamic content directly, which allows for a more flexible modeling of scene variations [6, 40, 8, 41].
-However, obtaining a high-quality representation with these methods often requires a large number of Gaussians, leading to substantial storage overhead.
-Furthermore, as shown in Figure 2, spatio-temporal methods face the challenge of effectively adjusting the Gaussian parameters when applied to dynamic scenes with rapid and complex motions, such as those in [18].
-This is alleviated by Dynamic 3D Gaussians [26], which enforces consistency across all frames, but at the cost of increased storage and rendering time.
+Dynamic scene reconstruction is essential for applications in virtual and augmented reality, gaming, and robotics, where a real-time and accurate representation of moving objects and their environment is the key to immersive experiences.
+Recent progress such as Neural Radiance Fields (NeRF) [mildenhall2021nerf] has enabled high-fidelity scene generation, although at the cost of a slow rendering speed.
+To address this limitation, 3D Gaussian Splatting (3DGS) [kerbl20233d] leverages sparse Gaussian splats for efficient scene rendering, particularly for static scenes.
+Since the advent of 3DGS, various methods have been proposed to extend dynamic scene modeling, either by evolving Gaussians over time to capture time-varying properties [4dgs, luiten2023dynamic, das2024neural, yang2024deformable, xu20234k4d] or by learning spatio-temporal Gaussians for more flexible scene representation [das2024neural, yang2024deformable, duan20244d, yang2023real, Jiang_2024_CVPR].
+Our experiments with such frameworks, illustrated in Figure 1 and in supplementary videos have highlighted that many state-of-the-art methods struggle to effectively adjust Gaussian parameters in dynamic scenes with rapid and complex motions, such as those of [Joo_2019].
+The only exception is Dynamic 3D Gaussians [luiten2023dynamic], which mitigates this issue by enforcing consistency across frames.
+However, this comes with a dramatic increase in storage and rendering costs, making this method impractical for real-world applications, such as deployment on VR/AR headsets.
 To address these challenges, we propose Temporally Compressed 3D Gaussian Splatting (TC3DGS), a novel approach designed to efficiently compress dynamic 3D Gaussian representations for high-quality, real-time scene rendering.
-Unlike traditional methods, TC3DGS reduces both the number and the memory footprint of the Gaussians, by selectively pruning splats based on temporal importance and learning a parameter-specific bit-precision.
+Unlike traditional methods, TC3DGS reduces both the number and the memory footprint of the Gaussians by selectively pruning splats based on temporal importance and learning a parameter-specific bit-precision.
 This selective compression allows us to maintain scene fidelity while significantly reducing storage and computational requirements, making TC3DGS well-suited for dynamic environments with complex motions.
-Our approach begins with a pruning and masking strategy designed to eliminate redundant Gaussians.
-While previous works [42, 11, 21] have proposed pruning methods for 3DGS, they do not attempt to model dynamic scenes and thus take no advantage of temporal compression.
-Unfortunately, adapting these pruning strategies to dynamic scenes is not straightforward, as different Gaussians may need to be pruned across different frames.
-Here, we introduce a method that explicitly integrates this temporal aspect into the training objective, allowing us to prune dynamic Gaussian splats more effectively.
-To further optimize memory usage in our scene representation, we also aim to compress the storage size of the remaining Gaussians.
-To this end, we developed a gradient-aware mixed-precision quantization method that adjusts the bit precision of each Gaussian parameter based on its sensitivity.
-We use gradient magnitudes to determine parameters with high sensitivity, and allocate them a higher precision, while those with lower impact on the scene are quantized with fewer bits.
-Therefore, our method achieves a good balance between compression and reconstruction accuracy.
-Finally, to further enhance the efficiency of our dynamic representation, we introduce a keypoint extraction algorithm as a post-processing step.
-This algorithm simplifies the temporal trajectories of the Gaussian parameters, preserving only the key data points that are crucial for accurate dynamic scene representation.
-This greatly reduces the amount of temporal data stored, allowing for more compact scene representations, as demonstrated in our experiments where we achieve a compression rate of up to 67$\times$ while preserving rendering quality.
-Our experiments on benchmark datasets demonstrate the benefits of our method across diverse scenarios.
+[…]
+Our experiments on benchmark datasets demonstrate the benefits of our method on diverse scenarios.
 Furthermore, we perform an ablation study to demonstrate the contribution of each key component in our pipeline.
 In a nutshell, our major contributions are as follows:
+  • We identify and analyze the failure of existing methods in handling scenes with rapid and complex motion, highlighting these shortcomings through qualitative and quantitative evaluations.
   • We introduce the first method to prune dynamic Gaussian splats, going beyond previous pruning techniques, designed for static scenes, by incorporating temporal relevance into the pruning process.
-  • We develop a sensitivity-driven, gradient-based quantization method that dynamically assigns bit precision to parameters based on their impact on reconstruction accuracy, optimizing memory usage.
+  • We develop a sensitivity-driven, gradient-based mixed-precision quantization method that dynamically assigns bit precision to parameters based on their impact on reconstruction accuracy, optimizing memory usage.
   • We propose a keypoint extraction post-processing algorithm to further reduce storage requirements by simplifying the Gaussian parameter trajectories, retaining only essential data points for compact scene representations.
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Ground Truth Dynamic 3DGS [26] TC3DGS (Ours) STG [23] 4D Gaussian [38]
-Figure 2: Comparative Evaluation on Panoptic Dataset [18] Recent lightweight state-of-the-art methods (STG [23], and 4D-Gaussian [38]) struggle to accurately reconstruct scenes in complex environments.
-In contrast, our compression strategy effectively captures and maintains high fidelity in dynamic scene details.
 2 Related Work
 Dynamic 3D reconstruction.
-Recent advancements in 3D reconstruction based on Neural Radiance Fields (NeRFs) [28] and 3D Gaussian Splatting (3DGS) [20] have achieved remarkable levels of visual fidelity and accuracy.
-These methods have subsequently been extended to 4D representations [33, 24, 13], enabling dynamic scene reconstruction.
-Methods to decompose a 4D scene into multiple 2D planes to learn a more compact representation are also explored by various methods [16, 3, 35, 1].
+Recent advancements in 3D reconstruction based on Neural Radiance Fields (NeRFs) [mildenhall2021nerf] and 3D Gaussian Splatting (3DGS) [kerbl20233d] have achieved remarkable levels of visual fidelity and accuracy.
+These methods have subsequently been extended to 4D representations [pumarola2021d, liu2023robust, TiNeuVox, kulhanek2024wildgaussians, Li2024ST], enabling dynamic scene reconstruction.
+Methods to decompose a 4D scene into multiple 2D planes to learn a more compact representation are also explored by various methods [kplanes_2023, Cao2023HexPlane, shao2023tensor4d, attal2023hyperreel].
 In the case of 3DGS, where Gaussians are explicitly stored and rendered, different approaches have emerged to model their time dependence.
-One prominent line of work [39, 26, 6, 40] in this area optimizes a canonical set of Gaussians from the initial frame, and combines it with a deformation motion field allowing temporal variations of the Gaussian parameters.
-However, these methods are limited to short videos, as they cannot add Gaussians after the initial frame.
-Another class of methods [41, 8, 19, 9] directly models temporal Gaussians that can be present in a subset of frames, enabling certain elements to appear in selected time ranges and thus increasing the expressivity of their reconstruction.
+One prominent line of work [4dgs, luiten2023dynamic, das2024neural, yang2024deformable, huang2024scgs, kratimenos2024dynmf] in this area optimizes a canonical set of Gaussians from the initial frame, and combines it with a deformation motion field allowing temporal variations of the Gaussian parameters.
+However, these methods are limited to short videos, as they cannot add Gaussians after the initial frame.Recent works [sun20243dgstream, hicom2024] have explored online causal training of 3DGS models for streaming applications and to handle long scenes.
+However, these methods scale with the number of frames, like Dynamic 3DGS [luiten2023dynamic].
+Another class of methods [yang2023real, duan20244d, katsumata2025compact] directly models temporal Gaussians that can be present in a subset of frames, enabling certain elements to appear in selected time ranges and thus increasing the expressivity of their reconstruction.
 However, a major limitation across these methods is that both training and inference times for novel view synthesis scale with the number of Gaussians, the length of the sequence, and the complexity of their parameters, presenting a key bottleneck in enhancing reconstruction quality.
@@
 An important research direction has thus emerged in developing more compact representations of radiance fields.
-For NeRFs, compact grid structures [30, 4, 15, 14] have already proven effective in reducing network sizes and enhancing accuracy.
-With 3DGS, recent works have either concentrated on optimizing the representation of the Gaussian parameters [11], or on identifying low-importance Gaussians and pruning them entirely [42, 11, 21].
-Unorthodox techniques like representing the Gaussian parameters as 2D grids and applying image compression techniques [29, 31] have also been studied.
-Better initialization and weighted sampling based pruning [12] has also shown promising results.
-Additionally, the use of traditional compression techniques such as vector quantization [21, 31] or entropy models [5] have shown some potential for the compression of static scenes, but scaling them to dynamic scenes with possibly hundreds or thousands of frames remains a challenge.
+For NeRFs, compact grid structures [muller2022instant, chen2022tensorf, fridovich2022plenoxels, franke2024trips, scaffoldgs] have already proven effective in reducing network sizes and enhancing accuracy.
+With 3DGS, recent works have either concentrated on optimizing the representation of the Gaussian parameters [fan2023lightgaussian, chen2024fast], or on identifying low-importance Gaussians and pruning them entirely [zhang2024lp, fan2023lightgaussian, lee2024compact, girish2024eaglesefficientaccelerated3d].
+Unorthodox techniques like representing the Gaussian parameters as 2D grids and applying image compression techniques [morgenstern2024self_grid, navaneet2024compgs] have also been studied.
+Better initialization and weighted sampling based pruning [Fang2024MiniSplattingRS] has also shown promising results.
+Additionally, the use of traditional compression techniques such as vector quantization [lee2024compact, navaneet2024compgs] or entropy models [chen2025hac] have shown some potential for the compression of static scenes, but scaling them to dynamic scenes with possibly hundreds or thousands of frames remains a challenge.
 Indeed, dynamic scenes require an even larger set of parameters to accurately capture motion, temporal variations, and complex interactions within the scene.
@@
 Our approach combines temporal relevance-based pruning, gradient-based mixed-precision quantization, and trajectory simplification to address the unique requirements of dynamic scenes.
-3 Method
-In this section, we will first briefly discuss the Dynamic 3DGS [26] method in Sec.
-3.1, which serves as the foundation of our proposed approach.
-We then detail in Sec.
-3.2.1 our novel masking and pruning strategies designed to eliminate redundant Gaussians.
-Next, we present a sensitivity-based mixed precision technique for efficient parameter compression in Sec.
-3.2.2, followed by a post-training compression strategy aimed at minimizing storage overhead in Sec.
-3.2.3.
+3 On the Challenges of Fast Motion
+One of the most widely used datasets for benchmarking dynamic novel view synthesis, the Neural 3D Video dataset [li2022neural3dvideosynthesis], consists exclusively of indoor scenes with objects moving within their local vicinity.
+Existing methods thus perform well, as object motion remains limited.
+However, when we applied these methods to the Panoptic Sports dataset [Joo_2019], their performance deteriorated significantly due to the presence of fast moving objects that travel longer distances.
+Below, we analyze the underlying reasons why these methods struggle on such complex datasets.
+Existing Gaussian Splatting methods for dynamic scenes often incorporate motion priors to regulate the movement of the Gaussian points.
+These priors determine the type of motion for which the method is more suitable.
+The key distinction between different approaches is how they integrate these motion priors.
+Specifically, deformation field-based methods, such as 4D-Gaussians [4dgs] and E-D3DGS [bae2024ed3dgs], implicitly encode motion priors within the structure of neural networks, yielding flexible but less interpretable motion representations.
+These methods fails in the presence of large sudden movements due to the spectral bias of neural networks towards simpler functions [rahaman2019spectral].
+As an alternative, STG [stg] models point trajectories as polynomial curves.
+However, this approach restricts the representation to 3rd degree polynomials, which limits the types of motion it can represent.
+While higher-degree polynomials enable the representation of more complex motion, they also introduce significant storage overhead.
+Deformable 3DGS [yang2024deformable] impose explicit constraints, e.g., local rigidity, to regulate motion in a more interpretable manner.
+However, the use of a neural network in Deformable 3DGS to model the deformation field inherits the limitations of 4D-Gaussians and E-D3DGS discussed earlier.
+Among the methods discussed above only Dynamic 3DGS [luiten2023dynamic], deformable field and explicit constraints based model, is trained causally, i.e., learning Gaussian deformations frame by frame, instead of optimizing the Gaussians across all time frames simultaneously, as done by the other methods.
+This causal training strategy requires storing and optimizing point positions in every frame, offering greater flexibility in motion modeling despite the local rigidity constraint, as the Gaussians are optimized independently for each frame.
+However, this comes at a high storage and rendering cost.
+Therefore, in section˜4, we propose a method to significantly compress Dynamic 3DGS while preserving its ability to capture complex motions.
+In Figure 1, we provide a visual comparison illustrating how existing methods struggle with scenes featuring rapid and complex motion.
+In addition, we provide videos in the supplementary material for a more extensive evaluation.
+4 Compression
+In this section, we introduce our TC3DGS approach, encompassing novel masking and pruning strategies designed to eliminate redundant Gaussians, a sensitivity-based mixed precision technique for efficient parameter compression, and a post-training compression strategy to minimize storage overhead.
+The overview of our method can be visualized in Figure 2.
 Figure 2: Overview of our Temporally Compressed 3D Gaussian Splatting for Dynamic Scenese (TC3DGS) method.
 Our approach involves a temporally consistent masking strategy to select relevant 3D Gaussians across frames.
-The masked Gaussians are then pruned and quantized using a gradient-based, parameter-aware bit-precision quantization scheme, which optimizes memory usage while preserving scene fidelity.
-3.1 Dynamic 3D Gaussians
-Dynamic 3DGS [26] models dynamic scenes by allowing the Gaussians to move and rotate over time while enforcing that they have persistent color, opacity, and scale.
-This approach reconstructs a dynamic 3D scene over time using a series of images taken from multiple cameras across different time steps, along with the cameras’ intrinsic and extrinsic parameters.
-The method sequentially reconstructs a dynamic 3D scene by initializing each time step from the previous one and performing test-time optimization without additional training data.
-This is unlike most other concurrent methods, which attempt to optimize all frames jointly.
-This conceptual difference allows Dynamic 3DGS [26] to enforce consistency in the reconstructed Gaussians across frames.
-Additionally, this method enables tracking objects throughout the dynamic scene, as each 3D Gaussian has a unique correspondence across frames.
-Each dynamic scene $\mathcal{S}$ is parameterized by a set of Dynamic 3D Gaussians, with certain parameters fixed for the entire sequence and others varying over time.
-Specifically, each Gaussian retains a fixed 3D scale $(s_{x},s_{y},s_{z})$, color $(r,g,b)$, opacity logit $o$, and background logit $bg$ as determined in the initial frame.
-In contrast, parameters such as the 3D center $(x_{t},y_{t},z_{t})$ and 3D rotation, expressed as a quaternion $(q_{w,t},q_{x,t},q_{y,t},q_{z,t})$, evolve over time.
-This formulation enables the Gaussians to represent consistent scene elements across frames while dynamically adjusting their positions and orientations.
-During training, this 3D scene representation is iteratively updated using a differentiable renderer to minimize the photometric error between the rendered and input images.
-Specifically, each Gaussian influences a point $p$ in physical 3D space according to the standard (unnormalized) Gaussian equation weighted by its opacity, i.e.,
-$f_{i,t}(p)$ $=\sigma({o_{i}})\exp\left(-\frac{1}{2}(p-\mu_{i,t})^{T}\Sigma_{i,t}^{-1}(p-\mu_{i,t})\right)\;,$
-[…]
+The masked Gaussians are then pruned and quantized using a gradient-based, parameter-aware bit-precision quantization scheme.
 4.1 TC3DGS
-Dynamic 3DGS [26] is particularly promising because it models the dynamic scene as movements of Gaussians under kinematic constraints w.r.t.
+Dynamic 3DGS [luiten2023dynamic] is particularly promising because it models the dynamic scene as movements of Gaussians under kinematic constraints w.r.t.
 the previous timestep.
@@
 However, by learning position and rotation at each timestep separately, the number of parameters increases linearly, resulting in large storage sizes, and thus limiting its applicability to high-fidelity and long-range scene modeling.
+Figure 3: Mask Consistency. (Left) Mask values of Gaussians trained independently for each time frame. (Right) Mask values trained with our loss.
 4.1.1 Gaussian Masking and Pruning
-Pruning techniques aimed at identifying and removing low-importance Gaussians have been successfully applied to static 3D Gaussian splatting [21, 32, 17, 11].
-However, for dynamic scenes, a temporally consistent importance measure is required to ensure that the pruned Gaussians remain insignificant throughout the scene’s duration.
-Various approaches to computing the importance of each Gaussian in static scenes using training images and camera positions have bee proposed [32, 11].
+Pruning techniques aimed at identifying and removing low-importance Gaussians have been successfully applied to static 3D Gaussian splatting [lee2024compact, Niedermayr_2024_CVPR, girish2024eaglesefficientaccelerated3d, fan2023lightgaussian].
+However, for dynamic scenes, a temporally consistent importance measure is required to ensure that the pruned Gaussians remain insignificant throughout the scene duration.
+Various approaches to computing the importance of each Gaussian in static scenes using training images and camera positions have bee proposed [Niedermayr_2024_CVPR, fan2023lightgaussian].
 These methods focus on the contribution of each Gaussian to the training views.
@@
 To prune Gaussians effectively from dynamic scenes, it is essential to maintain the contributions of high-importance Gaussians consistently high, while suppressing low-importance ones, thus enabling more effective pruning.
-Compact-3DGS [21] introduces a masking approach based on Gaussian volume and transparency.
-Gaussians with low opacity, minimal volume, or both are masked out since they have negligible impact on the rendered images.
+Compact-3DGS [lee2024compact] introduces a masking approach based on Gaussian volume and transparency.
+Gaussians with low opacity, minimal volume, or both are masked out because they have negligible impact on the rendered images.
 A mask parameter $m\in\mathbb{R}^{N}$ is learned to produce binary masks $M\in\{0,1\}^{N}$ using a straight-through estimator.
@@
 $\hat{s}_{n}$ $=M_{n}s_{n},\quad\hat{o}_{n}=M_{n}o_{n},$ (2)
-where $n$ represents the Gaussian index, $\epsilon$ denotes the masking threshold, $\operatorname*{sg}(\cdot)$ is the stop-gradient operator, and $\mathds{1}[\cdot]$ and $\sigma(\cdot)$ correspond to the indicator and sigmoid functions, respectively.
-Compact-3DGS [21] explains that this mask learns to remove Gaussians with low opacity and/or small volume.
+where $n$ represents the Gaussian index, $\epsilon$ denotes the masking threshold, $\operatorname*{sg}(\cdot)$ is the stop-gradient operator, $s_{n}\in\mathbb{R}^{3}$ represents scale of the gaussian, $o_{n}\in\mathbb{R}$ represents opacity of the gaussian, and $\mathds{1}[\cdot]$ and $\sigma(\cdot)$ correspond to the indicator and sigmoid functions.
+In [lee2024compact], it is explained that this mask learns to remove Gaussians with low opacity and/or small volume.
 However, as shown in Figure 3, this explanation does not hold in dynamic scenes.
 The value of $m_{n}$ can change significantly across time frames, even with fixed opacity $o$ and scale $s$ after the initial time frame.
-Figure 4: Mask Consistency. (Left) Mask values of Gaussians trained independently for each time frame. (Right) Mask values trained with our loss.
 In dynamic scenes, $m_{n}$ depends on the 2D projected area of the Gaussians in the training views and their transmittance, where transmittance $T_{i}$ for the $i^{th}$ Gaussian along a camera ray is defined as the Gaussian’s contribution to blending, i.e.,
@@
 Similarly, when a Gaussian moves toward or away from the training cameras, its projected area in the views changes, affecting $m_{n}$ accordingly.
-Since the values in $m$ strictly need to be high due to their role in rendering through Eq.
-1 and the photometric loss, we introduce an additional regularization to incentivize lower values.
+Since the values in $m$ strictly need to be high due to their role in rendering through eq.˜1 and the photometric loss, we introduce an additional regularization to incentivize lower values.
 We thus regularize $m$ by minimizing
@@
 The flexibility of this learned mask is one of its key advantages, as it can be optimized to exhibit desired properties via the use of additional constraints.
-A key motivation for ensuring consistency of $m_{n,t}$ across frames is to capture the global importance of the Gaussians during pruning.
-To achieve this, we learn $m_{n,t}$ for each timestamp and introduce a consistency loss function that encourages $m_{n,t}$ to remain close to $m_{n,t-1}$.
+A key motivation for ensuring consistency of $m_{n}$ across frames is to capture the global importance of the Gaussians during pruning.
+To achieve this, we replace $m_{n}$ with a time dependent $m_{n,t}$ defined for every timestamp separately and use it instead of $m_{n}$ in the rendering.
+We introduce a consistency loss function that encourages $m_{n,t}$ to remain close to $m_{n,t-1}$.
 Specifically, we define this mask consistency loss function as
@@
 Our approach uses gradient-based sensitivities to dynamically assign bit precision to each parameter, based on its influence on reconstruction accuracy.
-By leveraging this adaptive, in-optimization quantization, each parameter adjusts its quantization scale [10] in real-time, preserving detail in the reconstructed scene.
-We first calculate the mean sensitivity for each parameter based on the gradients [43, 32], reflecting each parameter’s impact on the image reconstruction performance.
-We introduce a sensitivity coefficient $S(\theta)$, which represents the responsiveness of image quality $Q$ to changes in a Gaussian parameter $\theta$.
-This coefficient is formulated as
-$S(\theta)$ $=\frac{1}{\sum_{k=1}^{K}N_{k}}\sum_{k=1}^{K}\big{|}\frac{\partial Q_{k}}{\partial\theta}\big{|}.$ (7)
-where $K$ is the total number of training images used for reconstruction, $N_{k}$ is the number of pixels in the $k^{th}$ image, and $Q_{m}$ denotes the cumulative pixel intensity across the RGB channels in image $k$, serving as a proxy for image quality.
-The coefficient $S(\theta)$ quantifies how sensitive $Q$ is to variations in $\theta$, measured by the gradient $\frac{\partial Q}{\partial\theta}$.
-A higher gradient indicates that small adjustments in $\theta$ yield large changes in image quality, suggesting a greater impact of $\theta$ on the reconstruction.
+By leveraging this adaptive, in-optimization quantization, each parameter adjusts its quantization scale [Esser2020LEARNED] in real-time, preserving detail in the reconstructed scene.
+We first calculate the mean sensitivity for each parameter based on the gradients [Zhang_2024, Niedermayr_2024_CVPR], reflecting each parameter’s contribution on the image reconstruction performance.
+We introduce a sensitivity coefficient, formulated as
+$\vskip-8.5359ptS(\theta)$ $=\frac{1}{\sum_{k=1}^{K}N_{k}}\sum_{k=1}^{K}\big|\frac{\partial Q_{k}}{\partial\theta}\big|.$ (7)
+where $K$ is the number of training images used for reconstruction, $N_{k}$ is the number of pixels in the $k^{th}$ image, and $Q_{k}$ denotes the cumulative pixel intensity across the RGB channels in image $k$.
+The coefficient $S(\theta)$ quantifies the average change in pixel intensity from training views when parameter $\theta$ is changed by a small amount.
 By using this impact-based metric, we can effectively rank the parameters by their importance on image fidelity.
@@
 This standardization ensures that all the coefficients fall within a consistent range.
-Using the normalized sensitivity $\gamma\in[0,1]$, we assign a dynamic bit precision $b$ within a range of bits $b\in[b_{\text{min}},b_{\text{max}}]$, where $b\in\mathbb{Z}$ is calculated as
-$b=\left\lfloor\gamma\cdot(b_{\text{max}}-b_{\text{min}})\right\rfloor+b_{\text{min}}.$ (8)
-Figure 5: Keypoint Interpolation.
-In this example, we represent a position across 150 frames with only 5, 4 and 6 keypoints for $x$, $y$ and $z$, respectively, with only $0.038$ MSE.
-By comparison, uniformly sampling 7 keypoints increases storage and increases error to $0.089$ MSE.
-This approach allocates higher bit precision to more sensitive parameters and lower bit precision to less sensitive ones, optimizing the balance between computational efficiency and model accuracy.
+Therefore, we allocate higher bit precision to more sensitive parameters and lower bit precision to less sensitive ones, optimizing the balance between computational efficiency and model accuracy.
 After running our scene reconstruction process for a specified number of iterations, we apply mixed-precision quantization to all Gaussian parameters, excluding the position parameter, to achieve low-bit precision for the other parameters.
-Instead of relying on traditional vector quantization (VQ) or basic min-max quantization, we propose a parameter quantization technique with learnable scaling factors [10], integrating it directly into the optimization process rather than treating it as a post-optimization fine-tuning step.
-Consequently, many parameters can be effectively quantized to even 4-bit precision, reducing memory and computational load without compromising reconstruction quality.
+Instead of relying on traditional vector quantization (VQ) or basic min-max quantization, we propose a parameter quantization technique with learnable scaling factors [Esser2020LEARNED], integrating it directly into the optimization process rather than treating it as a post-optimization fine-tuning step.
+Consequently, many parameters can be effectively quantized to 4-bit precision, reducing memory and computational load without compromising reconstruction quality.
 Training.
 We train the Gaussian parameters to model the scene one time frame at a time.
-Following [26], we use physically-based priors to regularize the Gaussians.
+Following [luiten2023dynamic], we use physically-based priors to regularize the Gaussians.
 The optimization objective is defined as
 $\mathcal{L}=\mathcal{L}_{original}+\lambda_{mask}\mathcal{L}_{mask}+\lambda_{mc}\mathcal{L}_{mc},$ (8)
-where $\mathcal{L}_{original}$ is the same loss function as in [26], while $\mathcal{L}_{mask}$ and $\mathcal{L}_{mc}$ are defined in Eq. 4 and 5, respectively.
-Algorithm 1 Keypoint Selection
+where $\mathcal{L}_{original}$ is the same loss function as in [luiten2023dynamic], while $\mathcal{L}_{mask}$ and $\mathcal{L}_{mc}$ are defined in Eq. 4 and 5, respectively.
+Algorithm 1: Keypoint Selection
 Input: values, max_keypoints ($max_{kp}$), tolerance ($\tau$)
@@
 5:  Compute error (mse) for entire sequence
-[…]
+6:  if mse $\leq\tau$ or $|\texttt{keypoints}|\geq max_{kp}$ then
 7:   break
@@
 4.1.3 Keypoint Interpolation
-Previous implementations of dynamic 3DGS often model the time-dependence of parameters by fitting polynomials to the Gaussian parameters, which greatly limits the range of motions that can accurately be modeled.
-By contrast, Dynamic 3DGS [26] opts for inefficiently storing all time-dependent parameters, such as Gaussian means, rotations, and colors, for all time frames.
+Dynamic 3DGS [luiten2023dynamic] opts for inefficiently storing all time-dependent parameters, such as Gaussian means, rotations, and colors, for all time frames.
 While this greatly increases the expressivity of this method compared to other works, it comes at a significant cost in memory.
+Figure 4: Keypoint Interpolation.
+In this example, we represent a position across 150 frames with only 5, 4 and 6 keypoints for $x$, $y$ and $z$, respectively, with only $0.038$ MSE.
+By comparison, uniformly sampling 7 keypoints increases storage and increases error to $0.089$ MSE.
 We take a different approach and observe that only a small subset of keypoints are required to accurately reconstruct complex motions.
+Instead of storing each gaussian’s position and rotation per frame, we identify keypoints where the trajectory of a gaussian changes significantly and interpolate linearly between them.
 However, the placement of these keypoints across time cannot be predetermined, as it depends on the individual Gaussians.
-For instance, background Gaussians require a single keypoint to cover the whole sequence, whereas moving objects will need substantially more keypoints.
-This motivates the development of our keypoint selection strategy, which we adapt from the Ramer–Douglas–Peucker (RDP) algorithm [7].
-It is applied as a post-processing step to further reduce storage requirements for the Gaussian parameters that change over time in dynamic scenes.
+For instance, background Gaussians require a single keypoint for the entire sequence, whereas moving objects will need substantially more keypoints.
+This motivates the development of our keypoint selection strategy, which we adapt from the Ramer–Douglas–Peucker (RDP) algorithm [RDP].
+It is applied as a post-processing step to further reduce storage requirements for the time-varying Gaussian parameters in dynamic scenes.
 While the RDP algorithm selects keypoints from a sequence based on a local error tolerance, $\xi$, we propose a novel keypoint selection method, as outlined in Algorithm 1.
-This method offers several advantages over the RDP algorithm.
-It provides greater flexibility by allowing control via both an acceptable tolerance value, $\tau$, and a maximum number of keypoints.
+It provides greater flexibility by allowing control via both an acceptable tolerance value, $\tau$, and a maximum number of keypoints $max_{kp}$.
 The parameter $\tau$ defines the maximum allowable Mean Squared Error (MSE) over the sequence.
 Unlike RDP, which is solely controlled by $\xi$, our method enables a hard maximum bound on the number of keypoints, allowing for more precise, fine-grained control.
+In Figure 4, we illustrate with an example how our method achieves lower MSE despite storing fewer keypoints.
 Following this, we flatten and transpose the time-dependent parameters from $\mathbb{R}^{T\times N\times D}$ to $\mathbb{R}^{ND\times T}$, transforming the data into $ND$ sequences of length $T$.
@@
 The sparsity of these matrices is controlled using the parameters $max_{kp}$ and $\tau$, and and we store these sparse matrices to minimize memory usage.
-Upon loading the scene data, we interpolate the keypoints across all timesteps to reconstruct the dense matrices, which are then reshaped to their original parameter dimensions.
 5 Experiments
-We compare our method with multiple techniques extending the 3D Gaussians approach to dynamic scenes.
-We evaluate the methods on diverse datasets covering different real-world and synthetic scenarios.
-Results on additional datasets are provided in the supplementary material along with more visualizations.
-4.1 Datasets
-4.1.1 Panoptic Sports Dataset
-We evaluate our method on the Panoptic Sports dataset, a subset from the Panoptic Studio dataset [18].
-This dataset contains 6 different scenes, each having 31 camera sequences spanning 150 frames.
-We use 4 cameras (0, 10, 15 and 30) for testing and the rest for training, following the convention set by [26].
-In addition to the images, we use the provided foreground/background segmentation to apply a segmentation loss to improve the results and prevent the background from moving.
-This dataset contains complex motions with objects moving quickly and over long distances.
-Table 1: Quantitative results on the Panoptic dataset. The best result is shown in bold, while the second-best result is underlined. The rendering resolution is set to 640$\times$460.
-Model PSNR (dB)↑ SSIM↑ LPIPS↓ FPS ↑ Storage (MB)↓
-Dynamic 3DGS [26] 28.7 0.91 0.17 760 1994
-STG [23] 20.5 0.78 0.44 429 19
-4D Gaussian [38] 27.2 0.91 0.18 40 62
-TC3DGS (Ours) 28.3 0.91 0.17 890 49
-4.1.2 Neural 3D Video Dataset
-The Neural 3D Video dataset [22] consists of 6 scenes with a number of cameras ranging from 18 to 21 and sequences of 300 frames.
-The images in this dataset are or resolution 2704$\times$2028, but we downsample them to 1352$\times$1014 for fair comparison with other methods.
-We hold out camera 0 for testing and do not apply any background loss, as background segmentations are not available for this dataset.
-4.2 Implementation Details
-For our experiments, we closely follow the hyperparameters outlined in Dynamic 3DGS [26].
-Specifically, for the compression strategy, we set both the mask weight parameter $\lambda_{\text{mask}}$ and the masking consistency loss parameter $\lambda_{\text{mask-cons}}$ to $0.01$ in our primary experiments.
-In terms of quantization, we use Gaussian parameters with bin sizes of $b_{\text{min}}=4$ and $b_{\text{max}}=8$, while positional data is quantized to 16-bit precision to ensure spatial accuracy.
-For the learnable quantization step size, we use a $0.02$ learning rate.
-Quantization is applied after 6000 iterations in the first scene for all experiments.
-For keypoint interpolation, we set the tolerance to $\tau=1\times 10^{-5}$ for the Panoptic dataset and $\tau=1\times 10^{-7}$ otherwise.
-Moreover, we restrict the maximum number of keypoints to $\text{max}_{\text{kp}}=30$ for the Panoptic dataset and $\text{max}_{\text{kp}}=60$ otherwise, which provides an effective balance between compression and temporal trajectory accuracy.
-Additional ablation studies with varying values of certain hyperparameters are provided in the supplementary material.
+We evaluate the methods on diverse datasets covering different real-world scenarios.
+Results on additional datasets, along with dataset and implementation details, are provided in the supplementary material, which also includes visualizations.
+Moreover, ablation studies comparing our quantization strategy to uniform quantization are also included.
+Figure 5: Our method lies on the Pareto frontier, achieving competitive performance with significantly smaller model size.
 5.1 Results
-Table 2: Quantitative comparisons with various competitive baselines on the Neural 3D Video dataset.
-represent methods trained with a 50-frame video sequence, requiring six models to complete the evaluation.“Storage” refers to the total model size for 300 frames.
-[…]
-4DGS [41] 31.57 96.69 3128.00MB
-4D Gaussian [38] 31.15 30.00 90MB
-C-D3DGS [19] 30.46 118.00 338.00MB
-Deformable 3DGS [40] 30.98 29.62 32.64MB
-E-D3DGS [2] 31.20 69.70 40.20MB
-STG [23] 32.04 273.47 175.35MB
-Dynamic 3DGS [26] 30.97 460.00 2772.00MB
-TC3DGS (Ours) 30.58 596.32 51.34MB
-(a) Ground-Truth
-(b) TC3DGS Render
-Figure 6: Qualitative Evaluation on the Neural 3D Video dataset.
-This figure shows a comparison between the ground truth and our reconstruction, demonstrating that our method achieves near-identical results (PSNR: 32.7) despite extreme compression.
-Table 3: Ablation study on the proposed contributions. ‘M’, ‘Q’ and ‘I’ denote masking, quantization, interpolation, respectively. ‘#Gauss’ means the number of Gaussians.
+As shown in Table 1 and Figure 1, our method achieves comparable results to Dynamic 3DGS [luiten2023dynamic], while using, on average, 40 times less storage.
+Similarly, the results on the Neural 3D Video dataset in Table 1 demonstrate competitive performance with a significantly smaller storage footprint and fastest rendering speed.
+These improvements are achieved without a substantial increase in training time compared to Dynamic 3DGS [luiten2023dynamic], as our method trains fewer Gaussians and just one additional parameter per Gaussian for masking.
+Method Panoptic Dataset Neural 3D Video Dataset
+PSNR ↑ FPS ↑ Storage ↓ PSNR ↑ FPS ↑ Storage ↓
+STG 20.5 429 19MB 32.04 273 175MB
+4D Gaussian 27.2 40 62MB 31.15 30 90MB
+E-D3DGS 17.3 31 130MB 31.20 69 40MB
+Deform. 3DGS 14.3 26 192.6MB 30.97 29 33MB
+Dynamic 3DGS 28.7 760 1994MB 31.38 460 2772MB
+HiCoM 26.6 456 71MB 31.17 247 270MB
+TC3DGS (Ours) 28.3 890 49MB 30.96 596 51MB
+Table 1: Quantitative comparisons on the Panoptic and Neural 3D Video datasets.
+The best result is shown in bold, and the second-best is underlined.
+STG on the Neural 3D Video dataset is trained on 50-frame sequences and requires six models for evaluation.
+On the Panoptic dataset, we were unable to obtain reasonable results for the STG baseline [stg], while 4D Gaussian [4dgs] produce very distorted images, as shown in Figure 1.
+This dataset demonstrates the strengths of explicit methods such as Dynamic 3DGS [luiten2023dynamic] and ours.
+While Dynamic 3DGS [luiten2023dynamic] effectively models complex motion and performs well on the Panoptic dataset, it is highly memory inefficient.
+In contrast, our model achieves similar performance with significantly lower storage.
+Figure 5 shows a comparison with other methods.
 […]
@@
 27.9 189K 299 MB 750 32.8 59 K 194 MB 583
-27.9 189K 46 MB 750 32.7 59 K 53 MB 583
-Refer to caption Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption Refer to caption
-Dynamic 3DGS [26] Pruned Quantized + Pruned Quantized + Pruned + KPI
-Figure 7: Ablation study. Adding the different components of our method comes at virtually no cost in visual accuracy but at significantly decreased storage and rendering cost.
-As shown in Table 1, our method achieves results comparable to Dynamic 3DGS [26], while utilizing, on average, 40 times less storage.
-Similarly, Table 2 presents the results of our method on the DyNeRF dataset, demonstrating competitive performance with a significantly smaller storage footprint.
-Furthermore, our approach delivers the fastest rendering speed.
-On the Panoptic dataset, we failed to obtain reasonable results for the STG baseline [23], while 4D Gaussian [38] produce very distorted images, as shown in Figure 2.
-This dataset demonstrates the strengths of explicit methods such as Dynamic 3DGS [26] and ours.
-While Dynamic 3DGS [26] is able to model the complex motion and perform well on the Panoptic dataset, it is highly memory inefficient.
-In comparison, our model performs equally well while using much less storage.
+27.9 189K 44 MB 750 32.7 59 K 53 MB 583
+Table 2: Ablation study on the proposed contributions. ‘M’, ‘Q’ and ‘I’ denote masking, quantization, interpolation, respectively. ‘#Gauss’ means the number of Gaussians.
 5.2 Ablation Studies
-We conduct an ablation study to evaluate the effectiveness of the individual components of our method.
-We report the results of this ablation in Table 3 on two scenes from the Panoptic dataset [18] and the Neural 3D Video dataset [22].
-These results show the importance of each component of our method in reducing the storage size of the 3DGS [26].
-In these experiments, our pruning strategy divides the number of Gaussians by 2 and by 5 folds for each scene.
-Then, sensitivity-aware quantization reduces the necessary storage space by 5, and keypoint interpolation by an additional 4 to 5 fold.
-Altogether, TC3DGS yields a compression ratio of 49 and 64, respectively.
-Most notably, this drastic size reduction comes at little to no cost in novel view synthesis, with an insignificant drop in PSNR of at most 0.4, which is barely perceptible.
-5 Limitations
-Our approach has certain limitations rooted in the nature of dynamic 3D Gaussian splatting (3DGS) and the constraints of compressing temporal data for dynamic scenes.
-First, unlike compression approaches for static 3DGS, we are limited in our ability to aggressively compress position parameters, as doing so would compromise temporal consistency and spatial accuracy in the dynamic setting.
-Additionally, as our method builds on Dynamic 3DGS, it inherits its inability to accurately reconstruct new objects that enter the scene after the initial frame.
-This restricts its use in scenarios requiring complete adaptability to scene changes.
-Nonetheless, our method excels at handling scenes with complex and fast motions, maintaining high fidelity and rendering efficiency despite these constraints.
-6 Conclusion
-We introduced Temporally Compressed 3D Gaussian Splatting (TC3DGS), a novel framework designed to achieve memory-efficient and high-speed reconstruction of dynamic scenes.
-Our method achieves up to 67x compression and up to three times faster rendering while retaining high fidelity in complex motion scenarios.
-Through selective temporal pruning and gradient-based quantization, TC3DGS minimizes memory usage with minimal impact on visual quality.
-While our method is effective, limitations remain in aggressively compressing position data and handling new objects that enter the scene mid-sequence.
-Future work will focus on bridging the gap between spatio-temporal methods and storage-efficient dynamic scene reconstruction, aiming to extend adaptability and further reduce storage requirements for complex, evolving environments.
-[…]
-Supplementary Material
-7 Technicolor Dataset [34]
-To further demonstrate the effectiveness of our method, we evaluate our method on the Technicolor Dataset.
-The results in Table 4 show that our method provides competitive performance with a significantly smaller storage footprint.
-Figure 11 shows a comparison between the ground truth and our reconstruction, demonstrating that our method achieves near identical results.
-Table 4: Per-Scene Results on the Technicolor Dataset. Our method achieves performance comparable to Dynamic3DGS [26] while significantly reducing the model size.
-Method Avg. Birthday Fabien Painter Theater Train
-[…]
-STG [23] $33.60$ $32.09$ $35.70$ $36.44$ $30.99$ $32.58$
-D. 3DGS [26] $32.12$ $30.82$ $33.62$ $34.73$ $31.12$ $30.67$
-Ours $31.9$ $30.33$ $33.15$ $34.62$ $31.01$ $30.57$
-[…]
-STG [23] - - - - - -
-D. 3DGS [26] $0.88$ $0.90$ $0.88$ $0.87$ $0.86$ $0.88$
-Ours $0.85$ $0.88$ $0.86$ $0.81$ $0.84$ $0.86$
-[…]
-STG [23] $0.08$ $0.04$ $0.11$ $0.10$ $0.13$ $0.04$
-D. 3DGS [26] $0.13$ $0.07$ $0.16$ $0.16$ $0.14$ $0.11$
-Ours $0.14$ $0.09$ $0.15$ $0.17$ $0.16$ $0.13$
-Table 5: Per-scene quantitative comparisons on the Panoptic Sports Dataset [18].
-Method Avg. Basket. Juggle Boxes Softball Tennis Football
-[…]
-4DGaussians [38] $27.2$ $26.68$ $27.47$ $27.02$ $27.44$ $27.31$ $27.39$
-STG [23] $20.45$ $21.60$ $19.93$ $20.65$ $19.44$ $20.82$ $20.23$
-D. 3DGS [26] 28.70 28.22 29.48 29.46 28.43 28.11 28.49
-Ours $27.81$ $27.92$ $29.15$ $28.28$ $27.96$ $25.97$ $28.00$
-[…]
-4DGaussians [38] 0.91 $0.90$ 0.92 0.91 0.92 0.92 0.92
-STG [23] $0.79$ $0.78$ $0.78$ $0.79$ $0.79$ $0.78$ $0.78$
-D. 3DGS [26] 0.91 0.91 0.92 0.91 $0.91$ $0.91$ $0.91$
-Ours $0.89$ $0.89$ $0.90$ $0.89$ $0.89$ $0.89$ $0.89$
-[…]
-4DGaussians [38] $0.11$ 0.14 0.10 0.10 0.10 0.11 0.10
-STG [23] 0.10 0.14 0.10 0.10 0.10 0.11 0.10
-D. 3DGS [26] $0.17$ $0.18$ $0.15$ $0.17$ $0.19$ $0.18$ $0.17$
-Ours $0.20$ $0.20$ $0.19$ $0.19$ $0.20$ $0.20$ $0.20$
-SIZE(MB)$\downarrow$
-4DGaussians [38] $63$ $66$ $59$ $63$ $57$ $72$ $65$
-STG [23] 19 19 19 20 19 22 19
-D. 3DGS [26] $2008$ $2161$ $1935$ $2021$ $2021$ $1915$ $2000$
-Ours $49$ $46$ $32$ $44$ $41$ $48$ $34$
-8 Detailed Results for Panoptic and Neural 3D Video Dataset
-We present detailed results for the two datasets discussed in Sec.
-4.1 of our paper, summarized in Table 5 and 8.
-These results include various metrics evaluated for each scene in both datasets, offering a granular view of performance.
-Additionally, we provide supplementary visualizations for the Neural 3D Video dataset, which further demonstrate the effectiveness of our method in capturing dynamic scenes with high fidelity.
-Finally, we present an ablation study over the main elements of our method in Figure 8.
-9 Additional Experiments
-We conducted several experiments to better understand hyperparameters of our method.
-9.1 Masking and Mask Consistency
-We vary the weight of the mask loss $\lambda_{\text{mask}}$ and the masking consistency loss $\lambda_{\text{mask-cons}}$ on the first 50 frames of the Basketball scene from the Panoptic Sports dataset.
-The results, shown in Table 6, indicate that increasing the weight of these parameters leads to greater pruning of Gaussians but results in reduced image quality.
-We found $\lambda_{\text{mask}}=\lambda_{\text{mask-cons}}=0.01$ to be an effective balance between maintaining image quality and minimizing model size.
-(a) SSIM $\uparrow$
-Mask Consistency Loss
-Mask Loss 0 0.01 0.05 0.1
-0.005 0.9087 0.9105 0.9109 0.9068
-0.01 0.9126 0.9140 0.9113 0.9092
-0.05 0.9036 0.9006 0.8986 0.8970
-0.1 0.8938 0.8956 0.8949 0.8972
-(b) PSNR $\uparrow$
-Mask Consistency Loss
-Mask Loss 0 0.01 0.05 0.1
-0.005 28.0 27.8 28.2 26.8
-0.01 28.1 28.3 26.8 26.9
-0.05 27.2 26.6 26.0 26.0
-0.1 27.1 27.4 26.6 26.7
-(c) LPIPS $\downarrow$
-Mask Consistency Loss
-Mask Loss 0 0.01 0.05 0.1
-0.005 0.1797 0.1769 0.1781 0.1815
-0.01 0.1746 0.1761 0.1748 0.1829
-0.05 0.2046 0.2090 0.2139 0.2172
-0.1 0.2345 0.2294 0.2344 0.2335
-(d) Number of Gaussians $\downarrow$
-Mask Consistency Loss
-Mask Loss 0 0.01 0.05 0.1
-0.005 208403 155720 143112 140897
-0.01 153757 112909 98017 93568
-0.05 49770 47110 29629 26154
-0.1 28972 31067 19472 15656
-Table 6: Impact of Mask Loss and Mask Consistency Loss.
-Increasing the weight of these masks leads to more aggressive pruning of Gaussians, but at the cost of rendering quality.
-Setting $\lambda_{mask}=0.01$ and $\lambda_{mask-cons}=0.01$ provides a good trade-off between quality and storage efficiency.
-9.2 Quantization
+We conduct an ablation study to evaluate the effectiveness of each component in our method.
+Table 2 presents results on two scenes from the Panoptic dataset [Joo_2019] and Neural 3D Video dataset [li2022neural3dvideosynthesis], demonstrating how each step contributes to reducing storage.
+Our pruning strategy reduces the number of Gaussians by 2 and by 5 folds, followed by sensitivity-aware quantization, which compresses storage by 5 times, and keypoint interpolation, adding a further  5 times reduction.
+Overall, TC3DGS achieves compression ratios of 49 and 64, respectively, with minimal impact on novel view synthesis, which is barely perceptible.
+Additional ablations are provided in the Supplementary.
+5.3 Mixed-Precision Quantization
 We ran experiments with sensitivity-aware quantization with different bit-ranges by varying $b_{\text{min}}$ and $b_{\text{max}}$ as well as using uniform bitwidth for all paramaters.
@@
 Whereas, compared to uniform quantization using lower bitwidth, our average bitwidth is slightly higher while image quality is improved considerably.
+Table 3: Comparison of Sensitivity-Aware and Uniform Bit Quantization.
+Our sensitivity-aware [4,8]-bit quantization achieves PSNR comparable to 8-bit precision while providing compression similar to uniform 5-bit quantization.
 Bit-precision PSNR SSIM LPIPS Compression
@@
 Ours [3,8] 26.2 0.84 0.28 7.3x
-Table 7: Comparison of Sensitivity-Aware and Uniform Bit Quantization.
-Our sensitivity-aware [4,8]-bit quantization achieves PSNR comparable to 8-bit precision while providing compression similar to uniform 5-bit quantization.
+6 Conclusion
+We have introduced Temporal Compressed 3D Gaussian Splatting (TC3DGS), a novel framework designed to achieve memory-efficient and high-speed reconstruction of dynamic scenes.
+Our approach achieves up to 67x compression and up to three times faster rendering while maintaining high fidelity in complex motion scenarios.
+Through selective temporal pruning and gradient-based quantization, TC3DGS significantly reduces memory usage with minimal impact on visual quality.
+While our method is effective, limitations remain in aggressively compressing position data and handling new objects that enter the scene mid-sequence.
+Future work will focus on bridging the gap between spatio-temporal methods and storage-efficient dynamic scene reconstruction, improving adaptability and further reducing storage requirements for complex, evolving environments.
+Supplementary
+7 Implementation Details
+For our experiments, we closely follow the hyperparameters outlined in Dynamic 3DGS [luiten2023dynamic].
+Specifically, for the compression strategy, we set both the mask weight parameter $\lambda_{\text{mask}}$ and the masking consistency loss parameter $\lambda_{\text{mask-cons}}$ to $0.01$ in our primary experiments.
+In terms of quantization, we use Gaussian parameters with bin sizes of $b_{\text{min}}=4$ and $b_{\text{max}}=8$, while positional data is quantized to 16-bit precision to ensure spatial accuracy.
+For the learnable quantization step size, we use a $0.02$ learning rate.
+Quantization is applied after 6000 iterations in the first scene for all experiments.
+For keypoint interpolation, we set the tolerance to $\tau=1\times 10^{-5}$ for the Panoptic dataset and $\tau=1\times 10^{-7}$ otherwise.
+Moreover, we restrict the maximum number of keypoints to $\text{max}_{\text{kp}}=30$ for the Panoptic dataset and $\text{max}_{\text{kp}}=60$ otherwise, which provides an effective balance between compression and temporal trajectory accuracy.
+Additional ablation studies with varying values of hyperparameters are provided in the supplementary.
+8 Additional Ablation Experiments
+We conducted several experiments to better understand hyperparameters of our method as shown in Figure 6.
+Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption
+Dynamic 3DGS [luiten2023dynamic] Pruned Quantized + Pruned Quantized + Pruned + KPI
+Figure 6: Ablation Study on Panoptic Sports Dataset
+8.1 Masking and Mask Consistency
+We vary the weight of the mask loss $\lambda_{\text{mask}}$ and the masking consistency loss $\lambda_{\text{mask-cons}}$ on the first 50 frames of the Basketball scene from the Panoptic Sports dataset.
+The results, shown in Table 4, indicate that increasing the weight of these parameters leads to greater pruning of Gaussians but results in reduced image quality.
+We found $\lambda_{\text{mask}}=\lambda_{\text{mask-cons}}=0.01$ to be an effective balance between maintaining image quality and minimizing model size.
+(a) SSIM $\uparrow$
+Mask Consistency Loss
+Loss 0 0.01 0.05 0.1
+0.005 0.909 0.911 0.911 0.907
+0.01 0.913 0.914 0.911 0.909
+0.05 0.904 0.901 0.899 0.897
+0.1 0.894 0.896 0.895 0.897
+(b) PSNR $\uparrow$
+Mask Consistency Loss
+Loss 0 0.01 0.05 0.1
+0.005 28.0 27.8 28.2 26.8
+0.01 28.1 28.3 26.8 26.9
+0.05 27.2 26.6 26.0 26.0
+0.1 27.1 27.4 26.6 26.7
+(c) LPIPS $\downarrow$
+Mask Consistency Loss
+Loss 0 0.01 0.05 0.1
+0.005 0.180 0.177 0.178 0.182
+0.01 0.174 0.176 0.175 0.183
+0.05 0.205 0.209 0.214 0.217
+0.1 0.235 0.229 0.234 0.234
+(d) Number of Gaussians $\downarrow$
+Mask Consistency Loss
+Loss 0 0.01 0.05 0.1
+0.005 208k 156k 143k 141k
+0.01 154k 113k 98k 94k
+0.05 50k 47k 30k 26k
+0.1 29k 31k 19k 16k
+Table 4: Impact of Mask Loss and Mask Consistency Loss.
+Increasing the weight of these masks leads to more aggressive pruning of Gaussians, but at the cost of rendering quality.
+Setting $\lambda_{mask}=0.01$ and $\lambda_{mask-cons}=0.01$ provides a good trade-off between quality and storage efficiency.
+(a) Interpolation Error and PSNR vs Maximum Keypoints
+(b) Interpolation Error and SSIM vs Maximum Keypoints
+Figure 7: Impact of increasing Maximum Keypoints The figure illustrates the trend of interpolation error and PSNR/SSIM as we increase $max_{kp}$.
+It highlights that the rate of change in interpolation error and PSNR/SSIM diminishes as we increase $max_{kp}$
+(a) Mean number of Keypoints
+(b) Distribution of Keypoints
+Figure 8: Distribution of Keypoints Across Different Parameters.
+The figure illustrates the mean number of keypoints for Gaussian parameters.
+It highlights that only color parameters reach $max_{kp}$ saturation, while positions and rotations can be effectively estimated with fewer keypoints.
 8.2 Keypoint Interpolation
@@
 In the case of the Technicolor dataset, despite its long sequences, we matched the experimental setup by using 50 timesteps per scene and reduced $max_{kp}$ to 15.
-Refer to caption Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption Refer to caption
-Dynamic 3DGS [26] Pruned Quantized + Pruned Quantized + Pruned + KPI
-Figure 8: Ablation Study on Panoptic Sports Dataset
-Table 8: Per-scene quantitative comparisons on the Neural 3D Video Dataset [22].
+9 Detailed Results for Panoptic and Neural 3D Video Dataset
+We present detailed results for the two datasets discussed in Section 5.1 of our paper, summarized in Table 5 and 6.
+These results include various metrics evaluated for each scene in both datasets, offering a granular view of performance.
+Additionally, we provide supplementary visualizations for the Neural 3D Video dataset, which further demonstrate the effectiveness of our method in capturing dynamic scenes with high fidelity.
+Finally, we present an ablation study over the main elements of our method in Figure˜6.
+Table 5: Per-scene quantitative comparisons on the Panoptic Sports Dataset [Joo_2019].
+Method Avg. Basket. Juggle Boxes Softball Tennis Football
+[…]
+4DGaussians [4dgs] $27.2$ $26.68$ $27.47$ $27.02$ $27.44$ $27.31$ $27.39$
+STG [stg] $20.45$ $21.60$ $19.93$ $20.65$ $19.44$ $20.82$ $20.23$
+D. 3DGS [luiten2023dynamic] 28.70 28.22 29.48 29.46 28.43 28.11 $28.49$
+HiCom [hicom2024] $26.59$ $26.35$ $27.15$ $25.09$ $28.38$ $25.91$ 28.62
+Ours $27.81$ $28.02$ $28.65$ $28.28$ $27.96$ $25.97$ $28.00$
+[…]
+4DGaussians [4dgs] 0.91 $0.90$ 0.92 0.91 0.92 0.92 0.92
+STG [stg] $0.79$ $0.78$ $0.78$ $0.79$ $0.79$ $0.78$ $0.78$
+D. 3DGS [luiten2023dynamic] 0.91 0.91 0.92 0.91 $0.91$ $0.91$ $0.91$
+HiCoM [hicom2024] $0.87$ $0.82$ $0.88$ $0.86$ $0.90$ $0.86$ $0.91$
+Ours $0.89$ $0.89$ $0.90$ $0.89$ $0.89$ $0.89$ $0.89$
+[…]
+4DGaussians [4dgs] $63$ $66$ $59$ $63$ $57$ $72$ $65$
+STG [stg] 19 19 19 20 19 22 19
+D. 3DGS [luiten2023dynamic] $2008$ $2161$ $1935$ $2021$ $2021$ $1915$ $2000$
+HiCoM [hicom2024] $71$ $71$ $71$ $76$ $70$ $69$ $70$
+Ours $49$ $46$ $32$ $44$ $41$ $48$ $34$
+Table 6: Per-scene quantitative comparisons on the Neural 3D Video Dataset [li2022neural3dvideosynthesis].
 Some methods only report part of the scenes.
@@
 2 excludes the Coffee Martini scene.
-“-” denotes results that are unavailable in prior work.
+“-" denotes results that are unavailable in prior work.
+STG on the Neural 3D Video dataset is trained on 50-frame sequences and requires six models for evaluation.
 Method Avg. Coffee Martini Cook Spinach Cut Roasted Beef Flame Salmon Flame Steak Sear Steak
 […]
-Neural Volumes [25] 1 $22.80$ $22.80$
-LLFF [27] 1 $23.24$ $23.24$
-DyNeRF [22] 1 $29.58$ $29.58$
-HexPlane [3] 2 $31.71$ $32.04$ $32.55$ $29.47$ $32.08$ $32.39$
-NeRFPlayer [36] $30.69$ 31.53 $30.56$ $29.35$ $31.65$ $31.93$ $29.13$
-HyperReel [1] $31.10$ $28.37$ $32.30$ $32.92$ $28.26$ $32.20$ $32.57$
-K-Planes [16] $31.63$ $29.99$ $32.60$ $31.82$ $30.44$ $32.38$ $32.52$
-MixVoxels-L [37] $31.34$ $29.63$ $32.25$ $32.40$ $29.81$ $31.83$ $32.10$
-MixVoxels-X [37] $31.73$ $30.39$ $32.31$ $32.63$ 30.60 $32.10$ $32.33$
-4D Gaussians [38] $30.67$ $27.34$ $32.46$ $32.90$ $29.20$ $32.51$ $32.49$
-Dynamic 3DGS [26] $30.97$ $27.32$ $32.97$ $31.75$ $27.26$ $33.24$ $33.68$
-STG [23] 32.05 $28.61$ 33.18 $33.52$ $29.48$ 33.64 33.89
-Ours $30.58$ $27.42$ $31.92$ 33.61 $27.12$ $32.48$ $33.04$
+Neural Volumes [Lombardi2019] 1 $22.80$ - - - $22.80$ - -
+LLFF [mildenhall2019llff] 1 $23.24$ - - - $23.24$ - -
+DyNeRF [li2022neural3dvideosynthesis] 1 $29.58$ - - - $29.58$ - -
+HexPlane [Cao2023HexPlane] 2 $31.71$ - $32.04$ $32.55$ $29.47$ $32.08$ $32.39$
+NeRFPlayer [song2023nerfplayer] $30.69$ 31.53 $30.56$ $29.35$ $31.65$ $31.93$ $29.13$
+HyperReel [attal2023hyperreel] $31.10$ $28.37$ $32.30$ $32.92$ $28.26$ $32.20$ $32.57$
+K-Planes [kplanes_2023] $31.63$ $29.99$ $32.60$ $31.82$ $30.44$ $32.38$ $32.52$
+MixVoxels-L [Wang2023ICCV] $31.34$ $29.63$ $32.25$ $32.40$ $29.81$ $31.83$ $32.10$
+MixVoxels-X [Wang2023ICCV] $31.73$ $30.39$ $32.31$ $32.63$ 30.60 $32.10$ $32.33$
+4D Gaussians [4dgs] $30.67$ $27.34$ $32.46$ $32.90$ $29.20$ $32.51$ $32.49$
+Dynamic 3DGS [luiten2023dynamic] $30.97$ $27.32$ $32.97$ $31.75$ $27.26$ $33.24$ $33.68$
+STG [stg] 32.05 $28.61$ 33.18 $33.52$ $29.48$ 33.64 33.89
+Ours $30.96$ $27.42$ $32.72$ 33.61 $27.12$ $32.48$ $33.04$
 […]
-Neural Volumes [25] 1 $0.295$ $0.295$
-LLFF [27] 1 $0.235$ $0.235$
-DyNeRF [22] 1 $0.083$ $0.083$
-HexPlane [3] 2 $0.075$ $0.082$ $0.080$ $0.078$ $0.066$ $0.070$
-NeRFPlayer [36] $0.111$ $0.085$ $0.113$ $0.144$ $0.098$ $0.088$ $0.138$
-HyperReel [1] $0.096$ $0.127$ $0.089$ $0.084$ $0.136$ $0.078$ $0.077$
-MixVoxels-L [37] $0.096$ $0.106$ $0.099$ $0.088$ $0.116$ $0.088$ $0.080$
-MixVoxels-X [37] $0.064$ $0.081$ $0.062$ $0.057$ $0.078$ $0.051$ $0.053$
-Dynamic 3DGS [26] $0.082$ $0.122$ $0.077$ $0.087$ $0.09$ $0.072$ $0.068$
-STG [23] 0.044 0.069 0.037 0.036 0.063 0.029 0.030
+Neural Volumes [Lombardi2019] 1 $0.295$ - - - $0.295$ - -
+LLFF [mildenhall2019llff] 1 $0.235$ - - - $0.235$ - -
+DyNeRF [li2022neural3dvideosynthesis] 1 $0.083$ - - - $0.083$ - -
+HexPlane [Cao2023HexPlane] 2 $0.075$ - $0.082$ $0.080$ $0.078$ $0.066$ $0.070$
+NeRFPlayer [song2023nerfplayer] $0.111$ $0.085$ $0.113$ $0.144$ $0.098$ $0.088$ $0.138$
+HyperReel [attal2023hyperreel] $0.096$ $0.127$ $0.089$ $0.084$ $0.136$ $0.078$ $0.077$
+MixVoxels-L [Wang2023ICCV] $0.096$ $0.106$ $0.099$ $0.088$ $0.116$ $0.088$ $0.080$
+MixVoxels-X [Wang2023ICCV] $0.064$ $0.081$ $0.062$ $0.057$ $0.078$ $0.051$ $0.053$
+Dynamic 3DGS [luiten2023dynamic] $0.082$ $0.122$ $0.077$ $0.087$ $0.09$ $0.072$ $0.068$
+STG [stg] 0.044 0.069 0.037 0.036 0.063 0.029 0.030
 Ours $0.089$ $0.129$ $0.081$ $0.088$ $0.094$ $0.075$ $0.071$
-(a) Interpolation Error and PSNR vs Maximum Keypoints
-(b) Interpolation Error and SSIM vs Maximum Keypoints
-Figure 9: Impact of increasing Maximum Keypoints The figure illustrates the trend of interpolation error and PSNR/SSIM as we increase $max_{kp}$.
-It highlights that the rate of change in interpolation error and PSNR/SSIM diminishes as we increase $max_{kp}$
-(a) Mean number of Keypoints
-(b) Distribution of Keypoints
-Figure 10: Distribution of Keypoints Across Different Parameters.
-The figure illustrates the mean number of keypoints for Gaussian parameters.
-It highlights that only color parameters reach $max_{kp}$ saturation, while positions and rotations can be effectively estimated with fewer keypoints.
+10 Technicolor Dataset [technicolor]
+To further demonstrate the effectiveness of our method, we evaluate our method on the Technicolor Dataset.
+The results in Table 7 show that our method provides competitive performance with a significantly smaller storage footprint.
+Figure˜9 shows a comparison between the ground truth and our reconstruction, demonstrating that our method achieves near identical results.
+[…]
+Method Avg. Birthday Fabien Painter Theater Train
+[…]
+STG [stg] $33.60$ $32.09$ $35.70$ $36.44$ $30.99$ $32.58$
+D. 3DGS [luiten2023dynamic] $32.12$ $30.82$ $33.62$ $34.73$ $31.12$ $30.67$
+Ours $31.9$ $30.33$ $33.15$ $34.62$ $31.01$ $30.57$
+[…]
+STG [stg] - - - - - -
+D. 3DGS [luiten2023dynamic] $0.88$ $0.90$ $0.88$ $0.87$ $0.86$ $0.88$
+Ours $0.85$ $0.88$ $0.86$ $0.81$ $0.84$ $0.86$
+[…]
+STG [stg] $0.08$ $0.04$ $0.11$ $0.10$ $0.13$ $0.04$
+D. 3DGS [luiten2023dynamic] $0.13$ $0.07$ $0.16$ $0.16$ $0.14$ $0.11$
+Ours $0.14$ $0.09$ $0.15$ $0.17$ $0.16$ $0.13$
+11 Limitations
+Our approach has certain limitations rooted in the nature of dynamic 3D Gaussian splatting (3DGS) and the constraints of compressing temporal data for dynamic scenes.
+First, unlike compression approaches for static 3DGS, we are limited in our ability to aggressively compress position parameters, as doing so would compromise temporal consistency and spatial accuracy in the dynamic setting.
+Additionally, as our method builds on Dynamic 3DGS, it inherits its inability to accurately reconstruct new objects that enter the scene after the initial frame.
+This restricts its use in scenarios requiring complete adaptability to scene changes.
+Nonetheless, our method excels at handling scenes with complex and fast motions, maintaining high fidelity and rendering efficiency despite these constraints.
 (a) Ground-Truth
@@
 This figure shows a comparison between the ground truth and our reconstruction, demonstrating that our method achieves near-identical results despite extreme compression.
-Top: Cook Spinach, Bottom: Sear Steak
+Top: Cook Spinach, Bottom: Sear Steak.
+11.1 Datasets
+11.1.1 Panoptic Sports Dataset
+We evaluate our method on the Panoptic Sports dataset, a subset of the Panoptic Studio dataset [Joo_2019].
+This dataset contains 6 different scenes, each having 31 camera sequences spanning 150 frames.
+We use 4 cameras (0, 10, 15 and 30) for testing and the rest for training, following the convention set by [luiten2023dynamic].
+In addition to the images, we use the provided foreground/background masks to apply a segmentation loss to improve the results and prevent the background from moving.
+This dataset contains complex motions with objects moving quickly and over long distances.
+11.1.2 Neural 3D Video Dataset
+[…]